### v1

# Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems

Abstract
Ensuring LLM reliability requires not only solving complex problems but also recognizing when a problem is unsolvable. Current models often struggle to distinguish objective unsolvability (inherent contradictions in the problem) from subjective capability limitations (problems beyond the model’s competence), which leads to hallucinations and overconfidence. To address this, we propose UnsolvableQA and UnsolvableRL to solve feasible problems, detect inherent contradictions, and prudently refuse tasks beyond capability. Specifically, we construct UnsolvableQA, a dataset of paired solvable and unsolvable instances derived via a dual-track methodology: programmatic generation for logic puzzles and a novel “Reverse Construction” method that injects contradictions into valid reasoning chains for mathematics. Building on this dataset, we introduce UnsolvableRL, a reinforcement learning framework with three reward components jointly accounting for accuracy, unsolvability, and difficulty. Empirical results show that our approach achieves near-perfect unsolvability detection while also improving accuracy on solvable tasks. Crucially, we identify Capability Collapse, demonstrating that explicit exposure to unsolvable data is indispensable for preventing models from becoming systematically overconfident. Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems
Dengyun Peng1,3thanks: Equal contribution   Qiguang Chen [^1] footnotemark: 1   Bofei Liu1   Jiannan Guan1   Libo Qin2thanks: Corresponding Author Zheng Yan1Jinhao Liu1Jianshu Zhang3Wanxiang Che122footnotemark: 2 1LARG, Research Center for Social Computing and Interactive Robotics, HIT 2School of Computer Science and Engineering, Central South University 3iFLYTEK
1 Introduction
As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably. Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning) and selective classification (fisch2022calibrated) during reasoning underscores the need for models to know when not to answer. Further, studies on model capability boundaries (chen2024unlocking) examine model self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware). More recently, TruthRL (wei2025truthrl) encourages models to refuse when uncertain through a ternary reward mechanism.
Figure 1: Comparison of alignment paradigms. We treat the solution space (solvable vs. unsolvable) and question difficulty (easy vs. hard) as orthogonal dimensions, requiring models to objectively detect contradictions and subjectively calibrate confidence.
However, a key limitation of these approaches is that they mainly focus on refusing questions that are solvable but currently beyond the model’s capabilities (“Need to Calibrate” region in Figure 1 (a)). This focus neglects a complementary and arguably more fundamental challenge: identifying problems that are inherently unsolvable due to logical contradictions. As illustrated by the gray “Be ignored” region in Figure 1 (a), existing models treat such an unsolvability as mere high difficulty or hallucinate answers even for simple contradictions.
To address this, our work distinguishes these scenarios by treating Solution Space (solvable vs. unsolvable) and Question Difficulty (easy vs. hard) as orthogonal dimensions (Figure 1 (b)). We argue that a reliable AI system must exhibit two behaviors: (1) Objective Detection, accurately identifying and flagging internally contradictory problems as “unsolvable” (top-left yellow quadrant); and (2) Subjective Calibration, safely refusing theoretically solvable problems that exceed the model’s current reasoning capacity. Without this distinction, models may confidently pursue hallucinations or unnecessarily reject solvable tasks.
First, to enable deep analysis of unsolvability, we present a fully automated pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning logic, puzzles, and mathematics. For programmatically verifiable tasks, we create provably unsolvable instances by systematically injecting logical contradictions at the code level. For domains such as mathematics that lack direct programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” methodology. This LLM-driven approach starts from a solvable problem and its correct Chain-of-Thought (CoT), inserts a controlled logical contradiction into the CoT, and reverse-engineers a corresponding unsolvable problem aligned with the flawed reasoning.
To train models that remain robust on both solvable and unsolvable inputs, we design UnsolvableRL, a dynamic reward alignment scheme based on GRPO (shao2024deepseekmath) that enables the model to detect inherent unsolvability and calibrate its confidence. Specifically, we assign distinct optimization objectives to two refusal types: (1) inherent unsolvability is formulated as a supervised classification task, whereas (2) capability-based calibration is governed by a dynamic confidence threshold with the model’s evolving performance.
Empirical results show a synergistic effect: UnsolvableRL achieves near-perfect precision in unsolvability detection and substantially higher reasoning accuracy on solvable tasks, nearly doubling the combined score. Our analysis also reveals a capability collapse phenomenon: without explicit exposure to unsolvable data, especially smaller models face a severe decline in self-monitoring, becoming overconfident in their outputs.
Our contributions are summarized as follows:
  • We propose a comprehensive alignment framework that treats objective unsolvability and subjective capability as orthogonal dimensions, establishing a robust three-way decision boundary for reliable reasoning.
  • We introduce UnsolvableQA, a dataset constructed via a novel “Reverse Construction” methodology. This approach bridges the data gap by systematically injecting logical contradictions into valid Chain-of-Thought traces to generate verifiable unsolvable instances.
  • We propose UnsolvableRL, a training paradigm leveraging dynamic confidence thresholds. This mechanism enables models to simultaneously learn to solve feasible tasks, detect inherent contradictions, and calibrate confidence on difficult problems.
  • Empirical results demonstrate synergistic improvements in both detection and reasoning. Crucially, we identify the phenomenon of Capability Collapse, revealing that explicit unsolvable data is essential to prevent model overconfidence.
2 Related Work
There are two primary areas related to our work: automated dataset construction for reasoning and reinforcement learning approaches for incentivizing selective refusal.
As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data). In the realm of reasoning, recent efforts primarily focus on ensuring data solvability and verifiability through formal or programmatic means. Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively. Concurrently, research on verifying the correctness of Chain-of-Thought (CoT) traces has also been advancing (ling2023deductive; chen2025towards). In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow. Although earlier benchmarks like SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) attempted to include unanswerable questions, their task formats are overly simplistic (primarily extractive span selection) and fail to effectively evaluate the refusal capabilities of modern models in complex logical reasoning scenarios. Consequently, existing data synthesis efforts remain largely concentrated on solvable instances, leaving a significant research gap.
Figure 2: The UnsolvableQA data collection pipeline. (a) Mathematical Data Collection: We employ a “Reverse Construction” method where contradictions are injected into the solution path of solvable problems. The resulting instances are verified by an LLM to ensure they lead to a contradiction (unsolvable) rather than a valid solution. (b) Puzzle Data Collection: We utilize domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
Beyond dataset construction, Reinforcement Learning (RL) (ouyang2022traininglanguagemodelsfollow; schulman2017proximalpolicyoptimizationalgorithms; rafailov2024directpreferenceoptimizationlanguage) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness. Recent RL algorithms for reasoning, such as GRPO (shao2024deepseekmath) and DAPO (yu2025dapoopensourcellmreinforcement), further enhance models’ reasoning abilities (sheng2025hybridflow). Concurrent work also leverages RL to improve reasoning efficiency under cost constraints (tu2025learning; team2025kimi; guo2025deepseek; wang2025aspoasymmetricimportancesampling; chen2025aware). In parallel, advances such as TruthRL (wei2025truthrl) introduce objectives that encourage models to refuse likely hallucinations, and research on self-awareness enables models to decline answering when appropriate (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; Steyvers2025).
Despite these advancements, existing frameworks largely conflate objective unsolvability with subjective uncertainty. By focusing primarily on solvable instances or confidence calibration, they overlook the necessity of explicitly learning to identify logical contradictions. We bridge this gap by systematically constructing guaranteed-unsolvable data to enforce a robust three-way decision boundary: solving feasible tasks, detecting inherent unsolvability, and Calibrating Refusal for problems beyond the model’s current capabilities.
3 UnsolvableQA Construction
Existing benchmarks mainly target solvable tasks, leaving models prone to hallucinations when judging feasibility. To address this issue, we introduce UnsolvableQA, a dataset for detecting problem inherent contradictions.
To achieve this, we use a dual-track methodology (illustrated in Figure 2) tailored to the specific verification constraints of different domains.
3.1 Mathematics Problem Generation via Reverse Construction
For mathematical reasoning tasks (e.g., algebra, geometry, combinatorics from AIME datasets), direct programmatic verification is often infeasible. We introduce a Reverse Construction methodology (see Figure 2 (a)) that leverages large language models to generate unsolvable problems by injecting logical contradictions into solvable seed problems, yielding high-quality unsolvable mathematics problems for training unsolvable policies.
Our reverse construction pipeline operates through two sequential stages:
Data Generation Process: First, we validate each seed problem by confirming that a strong reasoning model (DeepSeek-Reasoner) can correctly solve it and produce a complete chain-of-thought, ensuring sufficient reasoning capability. Then, we employ DeepSeek-Reasoner to design logical contradictions via two complementary strategies: Constraint Contradiction, which introduces conflicting external conditions, and Axiom Contradiction, which modifies reasoning steps to violate fundamental mathematical axioms. The model generates a structured plan specifying the location and mechanism of contradiction injection.
Validation Process: Based on this plan, the model then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction. Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-Reasoner to autonomously detect contradictions; if this fails, we explicitly provide the contradiction plan to assess genuine unsolvability. Only instances passing verification are retained.
3.2 Logic Puzzle Generation
To acquire strictly unsolvable questions in logic puzzle scenarios, we implement automated generators for four representative logic domains, following the sampling and verification process in Figure 2 (b). Each domain uses a deterministic solver to verify solutions and confirm unsolvability.
Game24. Solvability is checked by exhaustively enumerating all binary expression trees with exact rational arithmetic (Python Fraction) to avoid floating-point error. Unsolvable instances are obtained by rejection sampling, retaining number sets for which the exhaustive search proves that no solution exists.
Hamiltonian Cycle/Path. We encode graph traversability as a Boolean Satisfiability problem and verify randomly generated graphs using SAT solvers (MiniSat backend via pysat). Unsolvable instances are constructed by injecting structural impossibilities, such as removing critical edges to create bottlenecks or disconnecting components.
Hitori. We formulate Hitori as a Constraint Satisfaction Problem and, using python-constraint, generate puzzles with unique solutions. For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
Maze Navigation. Base mazes are generated via randomized Depth-First Search (DFS) to ensure nontrivial structure. To obtain unsolvable instances, we analyze all valid paths in a solvable maze and place an obstacle at a critical junction shared by every path, then verify unsolvability using Breadth-First Search (BFS).
Split Domain Solvable Unsolvable Total
Train Game24 50 50 100
Train Hamiltonian Cycle 48 48 96
Train Hamiltonian Path 50 50 100
Train Hitori 50 50 100
Train Maze 100 59 159
Train AIME (1983–2023) 50 32 82
Test Game24 50 50 100
Test Hamiltonian Cycle 48 50 98
Test Hamiltonian Path 50 50 100
Test Hitori 50 50 100
Test Maze(Easy) 100 94 194
Test AIME (2024–2025) 60 47 107
Train Total 348 289 637
Test Total 358 341 699
Table 1: Unified vertical statistics for UnsolvableQA train/test partitions. Train Maze has no difficulty split; AIME year ranges differ. (Note: AIME year ranges: train 1983–2023; test 2024–2025.)
3.3 Dataset Statistics and Quality Control
We construct UnsolvableQA to have approximately balanced solvable and unsolvable instances per domain, with minor residual imbalances caused by generator yields and verification constraints. Table 1 therefore reports aggregated statistics across the train and test partitions of UnsolvableQA. The test set contains 699 problems (358 solvable, 341 unsolvable), and the train set contains 637 problems (348 solvable, 289 unsolvable). The train partition pools all Maze instances without difficulty splits, whereas the test partition preserves Maze Easy. The AIME partitions cover different year ranges. Test prompts are diversified and paraphrased, while train prompts follow templated formats for scalable generation. All instances pass automated quality-control checks: logic puzzles are verified with multiple SAT/CSP/graph solvers, maze and programmatic tasks with deterministic solvers, and mathematically constructed problems with multi-pass LLM-based verification.
4 UnsolvableRL Methodology
4.1 Problem Formulation
We formulate the reasoning task as generating a response $y$ for a problem instance $x$, sampled from a distribution $\mathcal{D}$. Each instance $x\in\mathcal{D}$ is either objectively solvable ($x\in\mathcal{S}$) or unsolvable ($x\in\mathcal{U}$), and is further categorized as subjectively easy ($x\in\mathcal{E}$) or hard ($x\in\mathcal{H}$).
Specifically, to satisfy three distinct behavioral objectives, our training framework optimizes a policy $\pi_{\theta}$ as follows. First, for solvable and easy instances within the model’s capabilities ($x\in\mathcal{S}\cap\mathcal{E}$), the objective is Reasoning Accuracy, requiring the model to produce correct reasoning chains and final answers. Second, for inherently unsolvable instances ($x\in\mathcal{U}\cap\mathcal{E}$), the objective is Unsolvability Detection: the model must explicitly recognize the inconsistency and output a dedicated tag (e.g., <unsolvable>) instead of hallucinating a solution, which we treat as an objective correctness task. Third, for hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
Figure 3: UnsolvableRL Reward Framework. We decompose the alignment into two orthogonal components: (a) Objective Detection, which incentivizes the model to distinguish between solvable and unsolvable queries. Crucially, we impose a penalty (-0.5) for falsely labeling solvable problems as unsolvable to safeguard the model’s reasoning drive. (b) Subjective Calibration, which dynamically regulates refusal behavior. By comparing the target threshold ($\tau$) with the model’s rollout accuracy ($\beta$), the mechanism encourages refusal on hard tasks (where $\tau>\beta$) and suppresses it on easy ones (where $\tau<\beta$) via the dynamic reward term $\lambda(\tau_{i}-\beta_{i})$.
4.2 Group-Relative Policy Optimization
To efficiently optimize the policy across heterogeneous objectives without training a separate value function, this work adopts Group Relative Policy Optimization (GRPO). For each input $x$, the policy samples a group of $G$ outputs $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{G}\}$. The optimization objective is defined as:
$\begin{split}\mathcal{J}(\theta)&=\mathbb{E}_{x,\mathcal{Y}}\bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big(r_{i}(\theta)A_{i},\\ &\qquad\text{clip}(r_{i}(\theta),1-\epsilon,1+\epsilon)A_{i}\Big)\bigg],\end{split}$ (1)
where $r_{i}(\theta)=\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}$ is the probability ratio, and $\epsilon$ is the clipping parameter. GRPO computes the advantage $A_{i}$ for each output $y_{i}$ by normalizing rewards within the group instead of using a critic model:
$A_{i}=\frac{R(y_{i}\mid x)-\mu_{\mathbf{R}}}{\sigma_{\mathbf{R}}+\epsilon},$ (2)
where $\mu_{\mathbf{R}}$ and $\sigma_{\mathbf{R}}$ are the mean and standard deviation of the rewards $\mathbf{R}=\{R(y_{j}\mid x)\}_{j=1}^{G}$ within the group. This group-based normalization can stabilize training due to reward-scale differences across solvable and unsolvable domains.
4.3 Reward Design
As illustrated in Figure 3, our framework explicitly distinguishes between three types of model responses: (1) Reward for Reasoning Accuracy (providing a solution), (2) Reward for Unsolvability Detection (identifying inherent contradictions), and (3) Reward for Capability Calibration (admitting capability limits). We design a composite reward function that applies distinct mechanisms to these behaviors to align objective correctness with subjective capability.
Reward for Reasoning Accuracy ($R_{\text{acc}}$).
This component evaluates the correctness of the model’s judgments on verifiably solvable instances. (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance. (2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward. This asymmetric scheme emphasizes reasoning capability by rewarding successful problem solving while not explicitly penalizing unsuccessful but genuine reasoning attempts on solvable tasks.
Reward for Unsolvability Detection ($R_{\text{detect}}$).
This component evaluates how accurately the model identifies inherently contradictory or ill-posed instances. (1) A Correct-Detect Response (+1+1) is assigned when the model explicitly outputs the canonical tag unsolvable for such an instance. (2) To prevent degenerate behavior, we impose a False Unsolvability Penalty ($\rho=-0.5$): if the model incorrectly declares a solvable problem unsolvable, it receives a negative reward.
We explicitly motivate this penalty value via a decision-theoretic analysis (formal proof in Appendix A.1). Consider a model with an initial weak reasoning capability $\epsilon$ on solvable tasks (e.g., $\epsilon\approx 0.1$) and a posterior belief $p$ that a given problem is unsolvable. The model decides between two actions based on their expected rewards:
(1) Attempting to Solve: The expected reward depends on the problem being solvable and the model answering correctly: $\mathbb{E}[\text{Attempt}]\approx(1-p)\cdot\epsilon$.
(2) Declaring Unsolvable: The expected reward is the weighted sum of a correct rejection and a false rejection penalty: $\mathbb{E}[\text{Reject}]=p\cdot 1+(1-p)\cdot\rho$.
The model will only attempt to reason if $\mathbb{E}[\text{Attempt}]>\mathbb{E}[\text{Reject}]$. Solving this inequality yields a critical threshold for the belief $p$:
$p<\frac{\epsilon-\rho}{1+\epsilon-\rho}$
If we set $\rho=0$ (no penalty), the threshold collapses to $p<0.09$, meaning the model refuses to answer unless it is over 91% sure the problem is solvable. This leads to a local optimum of universal rejection. By setting $\rho=-0.5$, we relax this threshold to $p<0.375$. This “Threshold Expansion” effectively encourages the model to attempt reasoning even under moderate uncertainty, preventing the collapse of reasoning capabilities.
(3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.
Game24 HamCycle HamPath Hitori Maze(E) AIME24-25 Overall
S U M S U M S U M S U M S U M S U M $\mathrm{S}$ $\mathrm{U}$ $\mathrm{M}$
Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 67.9 25.3 46.6 48.4 36.3 42.4
+ Solvable-Only 85.5 11.5 48.5 41.1 65.0 53.1 52.0 66.5 59.3 53.5 5.0 29.3 95.0 83.0 89.0 68.3 18.9 43.6 65.9 41.7 53.8
+ UnsolvableRL-Mid 84.5 98.5 91.5 36.5 95.5 66.0 44.0 98.0 71.0 47.0 72.5 59.8 95.7 94.9 95.3 67.3 52.7 60.0 62.5 85.4 73.9
+ Fixed-$\tau$ 85.5 99.0 92.3 43.8 99.5 71.6 59.5 99.0 79.3 48.5 98.0 73.3 96.6 99.5 98.0 69.6 61.4 65.5 67.3 92.7 80.0
+ UnsolvableRL-Final 92.0 99.0 95.5 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.0 98.9 97.5 69.2 85.0 77.1 69.5 90.9 80.3
Qwen3-1.7B-Instruct 63.0 21.0 42.0 22.9 22.5 22.7 13.0 26.5 19.8 6.0 27.0 16.5 3.6 91.0 47.3 38.3 37.5 37.9 24.5 37.6 31.0
+ Solvable-Only 78.5 1.0 39.8 23.4 2.0 12.7 13.5 2.5 8.0 4.5 1.5 3.0 5.7 0.0 2.9 41.7 2.1 21.9 27.9 1.5 14.7
+ UnsolvableRL-Mid 70.5 90.5 80.5 19.3 85.5 52.4 17.0 91.5 54.3 11.0 56.0 33.5 4.1 92.0 48.1 39.2 39.6 39.4 26.9 75.9 51.4
+ UnsolvableRL-Final 79.0 99.5 89.3 21.9 87.5 54.7 17.0 87.5 52.3 6.0 56.5 31.3 8.6 96.8 52.7 37.3 48.4 42.8 28.3 79.4 53.9
Table 2: Per-domain performance for Qwen3 checkpoints. Values are percentages. For each test domain we report solvable accuracy (S), unsolvable rejection rate (U), and their mean (M = (S+U)/2). Overall metrics are unweighted means across all 6 domains. Model names indicate the training configuration and stage: UnsolvableRL is our main method trained on paired solvable/unsolvable data; Solvable-Only is an ablation trained exclusively on solvable instances (Step 60 for 1.7B, Step 120 for 4B); Fixed-$\tau$ is an ablation using a fixed calibration threshold. Suffixes Mid and Final represent intermediate (Step 60 for 1.7B, Step 120 for 4B) and extended (Step 360 for 1.7B, Step 320 for 4B) trained models, respectively.
Reward for Subjective Calibration ($R_{\text{cal}}$).
When the model outputs the specific refusal token (e.g., beyond my capabilities), we do not apply a static score. Instead, we regulate refusal via a target accuracy threshold $\tau\in[0,1]$. During each RL rollout, we compute the empirical accuracy $\beta$ of the current group as:
$\beta=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[y_{i}\text{ is correct}],$ (3)
where correctness includes both solving solvable problems and identifying unsolvable ones. The calibration reward is then defined as:
$R_{\text{cal}}(y)=\lambda(\tau-\beta)\,\mathbf{1}[y\text{ is a refusal}].$ (4)
Here, $(\tau-\beta)$ quantifies the reliability deficit—the gap between the target and the model’s actual performance. This creates a dynamic incentive mechanism as shown in Figure 3(b): when the model’s reasoning capability on the current batch is poor ($\beta<\tau$), the term is positive, encouraging refusal; conversely, when performance is high ($\beta>\tau$), the reward becomes negative, suppressing refusal on easy questions.
To enforce robust calibration, we implement a progressive schedule for $\tau$, increasing it steadily throughout training to approach $1$. This curriculum balances exploration and caution: in early stages, a lower $\tau$ provides a lenient margin to encourage the exploration of reasoning paths; in later stages, as $\tau\to 1$, it enforces a strict standard, compelling the model to answer prudently.
We mathematically justify the necessity of this dynamic schedule by analyzing the general condition for refusal. The model should choose to refuse only if the reward for refusal exceeds the expected reward of answering:
$R_{\text{cal}}>\mathbb{E}[R_{\text{ans}}].$ (5)
Note that the expected reward for answering is always non-negative ($\mathbb{E}[R_{\text{ans}}]\geq 0$). If we were to use a fixed threshold (e.g., $\tau=0.5$), as the model’s reasoning capability improves, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$). Once this occurs, the refusal reward becomes negative ($R_{\text{cal}}<0$). In this regime, the inequality $R_{\text{cal}}>\mathbb{E}[R_{\text{ans}}]$ becomes impossible to satisfy, as a negative reward cannot exceed a non-negative one. Consequently, for well-performed batches, the model is incentivized to suppress refusal entirely to avoid penalties. Our progressive schedule prevents this collapse by ensuring $\tau$ grows dynamically to maintain $\tau\gtrsim\beta$, keeping the refusal incentive positive for uncertain queries.
The final per-trajectory reward is the sum of these components:
$R=R_{\text{acc}}+R_{\text{detect}}+R_{\text{cal}}.$ (6)
Game24 HamCycle HamPath Hitori Maze(E) AIME24-25 Overall
Checkpoint Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R
Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 46.6 0.00 46.6 42.4 6.79 49.2
+ Solvable-Only 48.5 0.00 48.5 53.1 20.57 73.7 59.3 7.50 66.8 29.3 1.50 30.8 89.0 0.00 89.0 43.6 0.00 43.6 53.8 4.93 58.7
+ UnsolvableRL-Mid 91.5 0.00 91.5 66.0 19.01 85.0 71.0 13.00 84.0 59.8 1.50 61.3 95.3 0.00 95.3 60.0 0.00 60.0 73.9 5.59 79.5
+ Fixed-$\tau$ 92.3 0.00 92.3 71.6 17.19 88.8 79.3 10.00 89.3 73.3 1.00 74.3 98.0 0.00 98.0 65.5 0.00 65.5 80.0 4.70 84.7
+ UnsolvableRL-Final 95.5 0.00 95.5 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.5 0.00 97.5 65.9 0.00 65.9 80.3 8.01 88.3
Qwen3-1.7B-Instruct 42.0 0.00 42.0 22.7 9.38 32.1 19.8 9.25 29.0 16.5 0.50 17.0 47.3 0.00 47.3 37.9 0.21 38.1 31.0 3.22 34.3
+ Solvable-Only 39.8 0.00 39.8 12.7 10.94 23.6 8.0 7.00 15.0 3.0 0.25 3.3 2.9 0.00 2.9 21.9 0.10 22.0 14.7 3.05 17.8
+ UnsolvableRL-Mid 80.5 0.00 80.5 52.4 21.35 73.8 54.3 13.00 67.3 33.5 2.00 35.5 48.1 0.00 48.1 39.4 2.08 41.5 51.4 6.41 57.8
+ UnsolvableRL-Final 89.3 0.25 89.6 54.7 23.44 78.1 52.3 18.00 70.3 31.3 1.75 33.0 52.7 0.00 52.7 42.8 1.88 44.7 53.9 7.55 61.4
Table 3: Aggregated domain-level metrics for Qwen3 checkpoints. Each domain reports: Correctness (mean of solvable and unsolvable correct rates), Rejection (mean rejection rate), and Combined = Correctness + Rejection. Overall columns are unweighted means across domains.
4.4 Answer Extraction and Verification
For solvable tasks, deterministic verifiers (graph and constraint solvers, arithmetic evaluators, and reference-answer matchers) are applied. For unsolvable tasks, verification instead checks for the canonical unsolvable marker. For logic puzzles, domain-specific verifiers check the extracted answer against problem constraints and ground-truth solutions. For AIME and other mathematics problems, “math_verify” is used for symbolic equivalence checking. Answers containing “beyond my capabilities” are treated as calibration and handled by the dynamic reward logic.
5 Experiments
We evaluated Qwen3-1.7B and Qwen3-4B (yang2025qwen3technicalreport) with think mode after training and testing on UnsolvableQA. For evaluation, we report average@8 on the AIME solvable and unsolvable partitions to account for higher variance, and average@4 on all other datasets.
Figure 4: Evolution of Unsolvable Accuracy during training. The curves represent the average detection accuracy across unsolvable instances from the evaluated domains. Models trained with UnsolvableQA (solid lines) steadily improve their ability to identify contradictions. In contrast, the “No Data” ablation (dashed lines) leads to stagnation or, specifically for the 1.7B model, a catastrophic collapse in unsolvability detection.
UnsolvableRL substantially improves both accuracy and calibration.
As shown in Table 3, fine-tuning on a mixture of solvable and unsolvable data yields a consistent increase in overall accuracy across all test sets for both models. Aggregated metrics indicate that Qwen3-4B’s combined performance score (C+R) increases from 49.2% to 88.3% under the UnsolvableRL-Final configuration. Crucially, UnsolvableRL significantly improves unsolvable detection capabilities: the mean rejection rate on unsolvable instances ($U_{\mathrm{mean}}$) surged from 36.3% to 90.9% for Qwen3-4B, and from 37.6% to 79.4% for Qwen3-1.7B. In complex reasoning domains such as Game24 and Hamiltonian Path, the model attains rejection rates above 94% on unsolvable instances while maintaining, and in some cases improving, performance on solvable ones; for example, solvable accuracy on Game24 improves from 81.0% to 92.0%. Together, these results demonstrate that our training methodology enables the models to reliably distinguish between solvable and unsolvable problems and to exhibit the desired behavior of either answering correctly or abstaining.
Training on solvable-only data is fundamentally insufficient for learning robust calibration.
We further conducted an ablation study by training models on a variant of our data with all unsolvable instances removed (Solvable-Only). We compare the Solvable-Only model with the full-training (UnsolvableRL) model at the same RL steps. As shown in Table 2, the Qwen3-1.7B-Solvable-Only model’s mean unsolvable rejection rate collapses to 1.5%, compared to 75.9% for the UnsolvableRL model. This catastrophic failure underscores the necessity of explicitly including unsolvable examples during training to prevent over-confidence.
Dynamic thresholds achieve a more favorable trade-off between performance and calibration.
To assess the effectiveness of the Reward for Subjective Calibration, we compare our main model with a Fixed-$\tau$ variant that uses a static low threshold. As shown in Table 3, the dynamic mechanism substantially increases refusal under uncertainty, raising the mean rejection rate on solvable instances from 4.70% (Fixed-$\tau$) to 8.01%. This behavior arises from the interaction between model capability and reward. As training progresses and the model’s reasoning capability improves, the empirical accuracy $\beta$ increases. Under Fixed-$\tau$, once performance exceeds the static threshold ($\beta>\tau$), the calibration reward term $\lambda(\tau-\beta)$ becomes negative, penalizing refusal and effectively forcing the model to hallucinate answers on difficult problems. In contrast, our dynamic mechanism adjusts $\tau$ based on $\beta$, maintaining a positive incentive for refusal only when necessary. We provide a formal proof of this Capability-Refusal Trade-off in Appendix A.2. Consequently, the dynamic approach achieves a better balance, maintaining strong accuracy (Overall $S$ of 69.5% vs. 67.3%) alongside safer behavior.
Training without unsolvable instances leads to the collapse of unsolvability detection.
To further assess the role of negative supervision, we analyze the training dynamics in Figure 4. We track the Unsolvable Accuracy, defined as the accuracy of detecting unsolvable instances averaged across our evaluation domains. Models trained with our paired data (solid lines) steadily improve at detecting inherent contradictions, whereas ablation models without unsolvable samples (dashed lines) fail to generalize. This gap is most pronounced in the 1.7B model, which exhibits a sharp capability collapse, with its accuracy on unsolvable problems approaching zero. We attribute this to Gradient Interference in smaller models, where optimization for answering solvable questions suppresses the features required for detection (see Appendix A.3 for a theoretical derivation). These results demonstrate that explicit exposure to unsolvable problems is crucial for preventing overconfidence.
6 Conclusion
In this paper, we enhance the reliability of large language models (LLMs) by introducing two novel perspectives: detecting inherently unsolvable problems and calibrating model behavior on problems beyond their capabilities. To this end, we construct UnsolvableQA using a novel Reverse Construction methodology and align models with UnsolvableRL, which incorporates a dynamic refusal mechanism. Our approach enables models to effectively refuse both unsolvable and highly uncertain queries without degrading their reasoning performance. We expect this work to advance the development of LLMs that understand not only how to solve problems, but also when not to attempt them.
7 Future Work
We plan to extend this work in three directions. First, we will construct diverse Out-of-Distribution (OOD) test sets to rigorously assess the generalization of refusal boundaries. Second, we aim to empirically investigate the root cause of Capability Collapse in smaller models, verifying our gradient interference hypothesis via feature space analysis. Finally, we will validate the scalability of our framework by applying it to a broader range of model architectures.
Appendix A Appendix
A.1 Theoretical Justification for False Unsolvability Penalty
In this section, we provide a theoretical derivation to demonstrate why the penalty $\rho=-0.5$ is critical for preventing reward hacking, specifically the phenomenon of “collapse to universal rejection.”
Problem Setup. Let $x$ be an input problem. Let $\mathcal{S}$ and $\mathcal{U}$ denote the sets of solvable and inherently unsolvable problems, respectively. Based on our dataset distribution, we assume balanced prior probabilities $P(x\in\mathcal{S})=P(x\in\mathcal{U})=0.5$. Let the model have a weak initial reasoning capability $\epsilon$ on solvable tasks (e.g., $\epsilon\approx 0.1$), meaning the probability of generating a correct solution given a solvable attempt is $P(\text{correct}\mid x\in\mathcal{S},\text{attempt})=\epsilon$. Conversely, identifying unsolvability is a binary decision: if the model predicts “unsolvable”, it receives a reward of +1+1 if $x\in\mathcal{U}$ and a penalty $\rho$ if $x\in\mathcal{S}$ (False Unsolvability).
Decision Analysis. Consider the model’s decision process for a given input $x$ where the model has a posterior belief $p=P(x\in\mathcal{U})$ that the problem is unsolvable. The model compares the expected rewards $\mathbb{E}[R]$ of two actions: Attempting to Solve versus Rejecting.
For the Attempt action, the model tries to generate a solution. The expected reward is the probability the problem is solvable multiplied by the success rate $\epsilon$:
$\mathbb{E}[R_{\text{attempt}}]=(1-p)\cdot\epsilon\cdot(+1)+p\cdot 0=(1-p)\epsilon.$ (7)
Note that attempting to solve an inherently unsolvable problem yields 0 reward (hallucination).
For the Reject action, the model outputs “unsolvable”. The expected reward is the probability the problem is truly unsolvable plus the penalty for a false rejection:
$\mathbb{E}[R_{\text{reject}}]=p\cdot(+1)+(1-p)\cdot\rho.$ (8)
The Attempt Threshold. The model will choose to Attempt only when $\mathbb{E}[R_{\text{attempt}}]>\mathbb{E}[R_{\text{reject}}]$. Substituting the values, we get $(1-p)\epsilon>p+(1-p)\rho$. Solving for $p$, we derive the critical condition for attempting reasoning:
$p<\frac{\epsilon-\rho}{1+\epsilon-\rho}.$ (9)
This inequality defines the Attempt Threshold $T$. The model attempts reasoning only if its suspicion of unsolvability $p$ is lower than $T$.
Case 1: No Penalty ($\rho=0$). Substituting $\epsilon=0.1$ and $\rho=0$, we get $T_{\text{zero}}=\frac{0.1}{1+0.1}\approx 0.09$. This implies the model will only attempt to solve if it is more than 91% sure the problem is solvable. Given the uncertainty in early training stages (where $p$ is often near 0.5), the condition $p<0.09$ is rarely met. The optimal strategy collapses to “Always Reject”, as it guarantees a risk-free average reward of 0.5 (since 50% of data is unsolvable) compared to the meager 0.05 expected from attempting.
Case 2: With Penalty ($\rho=-0.5$). Substituting $\epsilon=0.1$ and $\rho=-0.5$, we get $T_{\text{penalty}}=\frac{0.1-(-0.5)}{1+0.1-(-0.5)}=\frac{0.6}{1.6}=0.375$. The threshold expands significantly from 0.09 to 0.375 (a $4\times$ increase). The model becomes willing to attempt reasoning even when it has moderate uncertainty about the problem’s solvability. This negative constraint effectively counteracts the low initial reasoning accuracy, forcing the model to explore solvable paths rather than collapsing into the local optimum of universal rejection.
A.2 Theoretical Analysis of Capability-Refusal Dynamics
In this section, we derive the mathematical relationship between the model’s reasoning capability and its tendency to refuse, explaining why a dynamic threshold is essential for preventing hallucination as the model becomes stronger.
Decision Model. For a given solvable input $x$, let the model’s internal confidence (probability of generating the correct answer) be $p_{\text{correct}}$. The model chooses between two actions: Answer or Refuse. The expected reward for answering is:
$\mathbb{E}[R_{\text{answer}}]=p_{\text{correct}}\cdot(+1)+(1-p_{\text{correct}})\cdot 0=p_{\text{correct}}$ (10)
The expected reward for refusing is determined by the current batch accuracy $\beta$ and the target threshold $\tau$:
$\mathbb{E}[R_{\text{refuse}}]=\lambda(\tau-\beta)$ (11)
The model will choose to Refuse only if the expected reward of refusing exceeds that of answering:
$\lambda(\tau-\beta)>p_{\text{correct}}\quad\implies\quad p_{\text{correct}}<\lambda(\tau-\beta)$ (12)
Dynamics under Training. During the training process, the model’s reasoning capability improves, which leads to two simultaneous changes: 1. For a specific difficult problem, the model’s confidence $p_{\text{correct}}$ increases. 2. The overall batch accuracy $\beta$ increases as the model solves more easy/medium problems correctly.
Failure of Fixed Threshold (Fixed-$\tau$). If we fix the threshold $\tau$ to a constant value (e.g., $\tau=0.5$), as the model converges, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$). When $\beta>\tau$, the term $\lambda(\tau-\beta)$ becomes negative. Substituting this into Inequality 12, the condition becomes:
$p_{\text{correct}}<\text{Negative Value}$ (13)
Since probability $p_{\text{correct}}\in[0,1]$ is always non-negative, this condition is impossible to satisfy. Therefore, the model will never refuse, even if its confidence $p_{\text{correct}}$ is extremely low (e.g., 0.01). It is forced to hallucinate an answer to avoid the penalty of refusal. This explains the lower rejection rate and lower accuracy observed in the Fixed-$\tau$ experiments.
Success of Dynamic Threshold. By dynamically setting $\tau$ (e.g., slightly higher than the moving average of $\beta$), we ensure that $(\tau-\beta)$ remains positive or near-zero. This maintains a valid decision boundary on the right side of Inequality 12, allowing the model to rationally refuse when its confidence $p_{\text{correct}}$ is truly low, regardless of how strong the model becomes overall.
A.3 Theoretical Analysis of Capability Collapse in Small Models
In this section, we propose a theoretical hypothesis for the capability collapse observed in smaller models (e.g., 1.7B) under the “No Data” setting. We frame this phenomenon as a potential instance of Gradient Interference arising from shared feature representations.
Problem Definition. Let $x$ be an input problem. We define the model’s refusal policy via a scalar logit $h_{\theta}(x)=w^{\top}\phi(x)$, where $\phi(x)$ represents the features in the last hidden layer and $w$ is the weight vector of the refusal classification head. The probability of rejection is $P(\text{reject}|x)=\sigma(h_{\theta}(x))$.
Training Dynamics under “No Data”. In the ablation setting, the training set $\mathcal{D}_{\text{train}}$ contains only solvable problems ($x_{s}\in\mathcal{S}$). The objective is to maximize the probability of answering (accepting) these problems. The loss function for a solvable sample $x_{s}$ can be formulated as minimizing the negative log-likelihood of not rejecting:
$\mathcal{L}(\theta)=-\log(1-P(\text{reject}|x_{s}))=-\log(1-\sigma(w^{\top}\phi(x_{s})))$ (14)
Computing the gradient with respect to $w$, we obtain:
$\nabla_{w}\mathcal{L}=\frac{1}{1-\sigma}\cdot\frac{\partial(1-\sigma)}{\partial h}\cdot\phi(x_{s})=\sigma(w^{\top}\phi(x_{s}))\cdot\phi(x_{s})=P(\text{reject}|x_{s})\cdot\phi(x_{s})$ (15)
Assuming simplified SGD with learning rate $\alpha$, the weight update is:
$\Delta w\propto-\nabla_{w}\mathcal{L}=-\alpha\cdot P(\text{reject}|x_{s})\cdot\phi(x_{s})$ (16)
Intuitively, whenever the model assigns a non-zero probability to rejecting a solvable problem (a false rejection), the optimizer pushes the weight vector $w$ in the direction opposite to $\phi(x_{s})$ to suppress this behavior.
Hypothesized Impact on Unsolvable Test Data. Consider an unsolvable test instance $x_{u}\in\mathcal{U}$. The resulting change in its refusal logit $h(x_{u})$ would be:
$\Delta h(x_{u})=(\Delta w)^{\top}\phi(x_{u})\propto-\alpha\cdot P(\text{reject}|x_{s})\cdot\left(\phi(x_{s})^{\top}\phi(x_{u})\right)$ (17)
This derivation suggests that the impact on unsolvable instances depends critically on the geometric relationship between the feature representations of solvable and unsolvable problems:
  • High Capacity Models (e.g., 4B+): We hypothesize that larger models can learn distinct, nearly orthogonal representations for solvable versus inherently contradictory problems (i.e., $\phi(x_{s})^{\top}\phi(x_{u})\approx 0$). In this case, suppressing rejection on $x_{s}$ has minimal impact on the model’s ability to reject $x_{u}$.
  • Low Capacity Models (e.g., 1.7B): Smaller models may suffer from feature entanglement, relying on shared heuristics (such as sentence complexity or topic) to represent both problem types. If $\phi(x_{s})$ and $\phi(x_{u})$ are positively correlated ($\phi(x_{s})^{\top}\phi(x_{u})>0$), the term $\Delta h(x_{u})$ becomes negative.
Conclusion. Under the assumption of feature entanglement in smaller models, training to answer solvable questions implicitly generates a negative gradient for the refusal logits of unsolvable questions. Without explicit negative samples ($x_{u}$) in the training set to provide a counter-gradient, this interference likely drives the refusal probability for unsolvable data toward zero, resulting in the observed capability collapse.
A.4 Logic Puzzle Generation Algorithms
A.4.1 Game24 Generation
The Game24 task involves combining a set of integers using arithmetic operations (+,,×,÷)(+,-$\times,\div$) to equal 24. To ensure absolute correctness, particularly for unsolvable instances, we employ an exhaustive search over all possible binary expression trees. Crucially, we utilize exact rational arithmetic (e.g., Python’s Fraction) rather than floating-point calculations to prevent precision errors that could misclassify instances (e.g., $23.99999\neq 24$).
For unsolvable instances, we use a rejection sampling strategy: we randomly sample integer sets and retain only those for which the exhaustive solver proves the solution set is empty.
Algorithm 1 Game24 Generation and Verification
0: Difficulty level $D$ (determines count $k\in\{4,5,6\}$), Target $T=24$
0: A paired instance $(S,\text{Label})$
1:Function Verify($S$):
2:  $\mathcal{T}\leftarrow$ All binary expression trees for set $S$
3:  for each tree $t\in\mathcal{T}$ do
4:   Evaluate $val\leftarrow t$ using Rational Arithmetic
5:   if $val==T$ then return Solvable, $t$
6:  end for
7:  return Unsolvable, $\emptyset$
8:End Function
9:repeat
10:  Sample a set of $k$ integers $S\leftarrow\{n_{1},\dots,n_{k}\}$ where $n_{i}\in[1,13]$
11:  $Status,Proof\leftarrow\textsc{Verify}(S)$
12:  if goal is Solvable and $Status==\textsc{Solvable}$:
13:   return $(S,Proof)$
14:  else if goal is Unsolvable and $Status==\textsc{Unsolvable}$:
15:   return $(S,\text{``Unsolvable''})$
16:until Valid instance found
A.4.2 Hamiltonian Cycle and Path Generation
The Hamiltonian Cycle problem asks whether a graph contains a closed loop visiting every vertex exactly once, while the Hamiltonian Path problem asks for a linear path visiting every vertex. Since these problems are NP-complete, we utilize a SAT-based verification pipeline to guarantee ground-truth labels.
SAT Encoding.
We encode the existence of a path/cycle as a Boolean Satisfiability Problem. Let $N$ be the number of vertices. We introduce boolean variables $x_{v,i}$ indicating that vertex $v$ occupies the $i$-th position in the sequence ($0\leq i<N$). The constraints are defined as follows: Bijection requires that $\sum_{v}x_{v,i}=1$ for all $i$ and $\sum_{i}x_{v,i}=1$ for all $v$. Adjacency mandates that for all $i\in[0,N-2]$, if $x_{u,i}$ and $x_{v,i+1}$ are true, then $(u,v)$ must be an edge in $E$. Finally, Cycle Closure (Cycle Only) ensures that if $x_{u,N-1}$ and $x_{v,0}$ are true, then $(u,v)\in E$. We use the Minisat solver (via pysat) to determine satisfiability.
Unsolvability Injection.
Instead of purely random sampling, we inject structural properties that inherently prevent Hamiltonian traversals. These include: Structural Bottlenecks, where graphs are generated with “bridge” nodes or articulation points that partition the graph into components that cannot be traversed in a single pass; Degree Constraints, where we inject nodes with degree <2<2 (dead ends) for cycles, or ensure the presence of >2>2 nodes with degree 1 for paths; and Disconnectivity, where we explicitly construct graphs with multiple connected components.
Algorithm 2 Hamiltonian Graph Generation
0: Type $T\in\{\text{Cycle},\text{Path}\}$, Node count NN
0: A graph $G=(V,E)$ with label
1:Function VerifySAT($G,T$):
2:  Construct CNF formula $\phi$ based on adjacency matrix of $G$
3:  Add bijection constraints to $\phi$
4:  if $T==\text{Cycle}$: Add closure constraint $v_{N-1}\leftrightarrow v_{0}$
5:  return SAT-Solver($\phi$)
6:End Function
7:repeat
8:  Initialize empty graph $G$ on NN nodes
9:  if goal is Solvable:
10:   Create a random permutation $P$ of vertices
11:   Add edges $(P_{i},P_{i+1})$ to form a base Path/Cycle
12:   Add random noise edges to increase density
13:  else if goal is Unsolvable:
14:   Select strategy $\sigma\in\{\text{Disconnect},\text{Bottleneck},\text{DeadEnd}\}$
15:   Construct $G$ satisfying $\sigma$ (e.g., partition $V$ into $V_{1},V_{2}$ with no edges between them)
16:   Add random noise edges strictly preserving $\sigma$
17:  $IsFeasible\leftarrow\textsc{VerifySAT}(G,T)$
18:until (IsFeasibleIsFeasible matches goal)
19:return $G$
A.4.3 Hitori Generation
Hitori is a logic puzzle played on an $N\times N$ grid. The goal is to shade cells such that no number appears more than once in any row or column, shaded cells do not touch orthogonally, and all unshaded cells remain connected.
We model the puzzle as a Constraint Satisfaction Problem (CSP). We define binary variables $x_{i,j}\in\{0,1\}$ for each cell ($1$ denotes shaded). The key constraints are: Uniqueness, which requires that if a number $v$ appears $k$ times in any row/column, at least k1k-1 occurrences must be shaded; Non-adjacency, ensuring that for any adjacent cells $u,v$, $x_{u}+x_{v}\leq 1$; and Connectivity, verified via BFS to ensure all cells where $x_{i,j}=0$ form a single connected component.
Solvable instances are filtered to ensure a unique solution. Unsolvable instances are those where the CSP solver (combined with connectivity checks) returns a solution count of 0.
Algorithm 3 Hitori Generation via CSP
0: Grid size NN
0: A validated Hitori grid $G$
1:Function SolveHitori($G$):
2:  Initialize CSP solver $P$
3:  Apply Row/Col constraints: Eliminate duplicates
4:  Apply Adjacency constraints: No adjacent black cells
5:  $\mathcal{S}_{local}\leftarrow P.\text{getSolutions()}$
6:  $\mathcal{S}_{valid}\leftarrow\emptyset$
7:  for each solution $s\in\mathcal{S}_{local}$ do
8:   if CheckConnectivity($s$) then $\mathcal{S}_{valid}.\text{add}(s)$
9:  end for
10:  return $|\mathcal{S}_{valid}|$
11:End Function
12:repeat
13:  Generate random $N\times N$ grid $G$ with values $\in[1,N]$
14:  $Count\leftarrow\textsc{SolveHitori}(G)$
15:  if goal is Solvable and $Count==1$:
16:   return $G$ {Unique solution found}
17:  else if goal is Unsolvable and $Count==0$:
18:   return $G$ {Provably no valid configuration}
19:until Valid instance found
A.4.4 Maze Navigation Generation
The Maze Navigation task requires finding a valid path from a start position (S) to an end position (E) in a grid-based maze, where movement is restricted by walls.
Verification Logic.
We employ Breadth-First Search (BFS) to determine solvability. BFS guarantees finding the shortest path if one exists. A maze is Solvable if a path from S to E exists, and Unsolvable if the set of reachable cells from S does not contain E.
Construction Pipeline.
We first generate a complex solvable maze using a randomized Depth-First Search (DFS) algorithm, which ensures a spanning tree structure. To increase complexity, we selectively remove walls to introduce cycles and multiple potential paths, while validating that the solution remains unique or manageable using BFS.
To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement. We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction—a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths). We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS. If the maze remains solvable (e.g., via a detour), we discard the instance or iteratively block remaining paths until connectivity is severed. This approach ensures that unsolvable mazes are structurally similar to solvable ones, often requiring the model to explore deep into the maze before realizing the path is blocked, thereby penalizing premature "give up" or hallucinated paths.
Algorithm 4 Maze Generation with Strategic Blockage
0: Dimensions $W,H$, Difficulty $D$
0: A maze grid $M$ and label $L$
1:Function Verify($M$):
2:  return BFS($M$, Start, End)
3:End Function
4:repeat
5:  $M_{base}\leftarrow$ GenerateSolvableDFS($W,H$)
6:  if goal is Solvable:
7:   return ($M_{base}$, Verify($M_{base}$))
8:  else if goal is Unsolvable:
9:   $\mathcal{P}\leftarrow$ All paths in $M_{base}$ from Start to End
10:   Select path $P\in\mathcal{P}$ based on difficulty $D$
11:   Select obstacle point $o\in P$ (excluding Start/End)
12:   $M_{blocked}\leftarrow M_{base}\cup\{o\}$
13:   if not Verify($M_{blocked}$):
14:    return ($M_{blocked}$, “Unsolvable”)
15:until Valid instance found
A.5 Hyperparameters
We implement our training pipeline using the verl (HybridFlow) framework (sheng2025hybridflow), which supports efficient reinforcement learning. The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 4. We set the group size $G$ (denoted as rollout.n in verl) to 12. Notably, we explicitly set use_kl_loss=False, meaning the KL divergence penalty term was disabled during optimization, relying instead on the probability ratio clipping to maintain policy stability. We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints.
Table 4: Hyperparameter settings used in our experiments. We employ Group Relative Policy Optimization (GRPO) with the KL divergence penalty explicitly disabled to encourage broader exploration within the solution space.
Hyperparameter Value
General Optimization & Generation
RL Algorithm GRPO
Group Size ($G$) 12
KL Penalty Term Disabled
PPO Mini-Batch Size 8
Sampling Temperature 0.6
Sampling Top-$k$ 45
Max Prompt Length 2048
Infrastructure
Hardware 8 $\times$ NVIDIA A100 GPUs
Tensor Parallelism (TP) 2
Model-Specific Settings
Qwen3-1.7B
   Training Batch Size $32\times 12$
   Max Response Length 32,000
Qwen3-4B
   Training Batch Size $14\times 12$
   Max Response Length 32,000

### v2

@@
 Abstract
-Ensuring LLM reliability requires not only solving complex problems but also recognizing when a problem is unsolvable.
-Current models often struggle to distinguish objective unsolvability (inherent contradictions in the problem) from subjective capability limitations (problems beyond the model’s competence), which leads to hallucinations and overconfidence.
-To address this, we propose UnsolvableQA and UnsolvableRL to solve feasible problems, detect inherent contradictions, and prudently refuse tasks beyond capability.
-Specifically, we construct UnsolvableQA, a dataset of paired solvable and unsolvable instances derived via a dual-track methodology: programmatic generation for logic puzzles and a novel “Reverse Construction” method that injects contradictions into valid reasoning chains for mathematics.
-Building on this dataset, we introduce UnsolvableRL, a reinforcement learning framework with three reward components jointly accounting for accuracy, unsolvability, and difficulty.
-Empirical results show that our approach achieves near-perfect unsolvability detection while also improving accuracy on solvable tasks.
-Crucially, we identify Capability Collapse, demonstrating that explicit exposure to unsolvable data is indispensable for preventing models from becoming systematically overconfident.
+Ensuring large language model (LLM) reliability requires distinguishing objective unsolvability (inherent contradictions) from subjective capability limitations (tasks exceeding model competence).
+Current LLMs often conflate these dimensions, leading to hallucinations in which they return confident answers to inherently unsolvable queries.
+To address this issue, we propose a multi-domain dataset containing both solvable and unsolvable questions, UnsolvableQA, together with an alignment framework, UnsolvableRL.
+First, we construct UnsolvableQA by “Reverse Construction” that systematically injects logical contradictions into otherwise valid reasoning chains.
+Second, we introduce UnsolvableRL, a reinforcement learning paradigm that balances objective unsolvability detection with calibrated confidence under capability limits.
+Empirically, our approach achieves near-perfect unsolvability detection ($90\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct.
+Crucially, we identify a data–training interaction: strict alignment constraints induce Capability Collapse without unsolvable data, but act as a regularizer for rigor when such data are included, thereby improving overall robustness.
 Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
@@
 As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably.
-Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning) and selective classification (fisch2022calibrated) during reasoning underscores the need for models to know when not to answer.
-Further, studies on model capability boundaries (chen2024unlocking) examine model self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware).
-More recently, TruthRL (wei2025truthrl) encourages models to refuse when uncertain through a ternary reward mechanism.
+Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning; fisch2022calibrated) during reasoning underscores the need for models to know when not to answer.
+Further, studies on LLMs’ capability boundaries (chen2024unlocking) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware).
+More recently, TruthRL (wei2025truthrl) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
 Figure 1: Comparison of alignment paradigms.
@@
 unsolvable) and question difficulty (easy vs.
-hard) as orthogonal dimensions, requiring models to objectively detect contradictions and subjectively calibrate confidence.
-However, a key limitation of these approaches is that they mainly focus on refusing questions that are solvable but currently beyond the model’s capabilities (“Need to Calibrate” region in Figure 1 (a)).
-This focus neglects a complementary and arguably more fundamental challenge: identifying problems that are inherently unsolvable due to logical contradictions.
-As illustrated by the gray “Be ignored” region in Figure 1 (a), existing models treat such an unsolvability as mere high difficulty or hallucinate answers even for simple contradictions.
-To address this, our work distinguishes these scenarios by treating Solution Space (solvable vs.
+hard) as orthogonal dimensions, requiring models to objectively detect solvability and subjectively calibrate confidence.
+However, a key limitation of these approaches is that they mainly focus on refusing questions that are solvable but currently beyond the model’s capabilities (the “Need to Calibrate” region in Figure 1 (a)).
+This emphasis overlooks a complementary, and more basic, challenge: identifying problems that are inherently unsolvable because they contain logical contradictions.
+More problematically, existing LLMs often misinterpret such unsolvability as high difficulty and may hallucinate answers even when the contradiction is explicit.
+To address this, our work treats Solution Space (solvable vs.
 unsolvable) and Question Difficulty (easy vs.
 hard) as orthogonal dimensions (Figure 1 (b)).
-We argue that a reliable AI system must exhibit two behaviors: (1) Objective Detection, accurately identifying and flagging internally contradictory problems as “unsolvable” (top-left yellow quadrant); and (2) Subjective Calibration, safely refusing theoretically solvable problems that exceed the model’s current reasoning capacity.
-Without this distinction, models may confidently pursue hallucinations or unnecessarily reject solvable tasks.
-First, to enable deep analysis of unsolvability, we present a fully automated pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning logic, puzzles, and mathematics.
-For programmatically verifiable tasks, we create provably unsolvable instances by systematically injecting logical contradictions at the code level.
-For domains such as mathematics that lack direct programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” methodology.
-This LLM-driven approach starts from a solvable problem and its correct Chain-of-Thought (CoT), inserts a controlled logical contradiction into the CoT, and reverse-engineers a corresponding unsolvable problem aligned with the flawed reasoning.
-To train models that remain robust on both solvable and unsolvable inputs, we design UnsolvableRL, a dynamic reward alignment scheme based on GRPO (shao2024deepseekmath) that enables the model to detect inherent unsolvability and calibrate its confidence.
-Specifically, we assign distinct optimization objectives to two refusal types: (1) inherent unsolvability is formulated as a supervised classification task, whereas (2) capability-based calibration is governed by a dynamic confidence threshold with the model’s evolving performance.
-Empirical results show a synergistic effect: UnsolvableRL achieves near-perfect precision in unsolvability detection and substantially higher reasoning accuracy on solvable tasks, nearly doubling the combined score.
-Our analysis also reveals a capability collapse phenomenon: without explicit exposure to unsolvable data, especially smaller models face a severe decline in self-monitoring, becoming overconfident in their outputs.
+Concretely, we argue that a reliable AI system must exhibit two behaviors: (1) Objective Detection, identifying internally contradictory problems and flagging them as “unsolvable” (yellow area); and (2) Subjective Calibration, safely refusing theoretically solvable but “difficult” problems that exceed the model’s reasoning capability.
+Without this distinction, models may hallucinate, forcing answers for inherently unsolvable problems.
+First, to support rigorous analysis of unsolvability, we develop a fully automated synthesis pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning puzzles and mathematics.
+For programmatically verifiable puzzle tasks, we construct provably unsolvable instances by injecting code-level logical contradictions.
+For mathematics problems without reliable programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” method: starting from a solvable problem and its correct rationale, we introduce a controlled contradiction into the rationale and then prompt LLMs to generate an unsolvable problem that matches the resulting flawed reasoning.
+Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (shao2024deepseekmath).
+Unlike prior approaches that conflate refusal types, we define 3 distinct RL objectives: (1) Inherent unsolvability rewards generating an explicit refusal tag when solvability is detected; (2) False unsolvability Penalty discourages erroneous refusals by penalizing labeling solvable problems as unsolvable, thereby requiring explicit solvability checks before refusal; (3) Capability calibration uses a mechanism that rewards refusal only when the problem difficulty exceeds the LLMs’ current capability.
+Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 40.3% to 90.7% (near-perfect) and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 80.1%.
+Our ablation study reveals three key insights: (1) models cannot learn to detect unsolvable problems without explicit unsolvable training data; (2) penalizing false refusals without such data causes “capability collapse”, where models stop refusing altogether; (3) when paired with unsolvable data, the penalty acts as a regularizer, trading a small drop in detection for stronger reasoning accuracy, a Pareto improvement over baselines.
 Our contributions are summarized as follows:
-  • We propose a comprehensive alignment framework that treats objective unsolvability and subjective capability as orthogonal dimensions, establishing a robust three-way decision boundary for reliable reasoning.
-  • We introduce UnsolvableQA, a dataset constructed via a novel “Reverse Construction” methodology.
-This approach bridges the data gap by systematically injecting logical contradictions into valid Chain-of-Thought traces to generate verifiable unsolvable instances.
-  • We propose UnsolvableRL, a training paradigm leveraging dynamic confidence thresholds.
-This mechanism enables models to simultaneously learn to solve feasible tasks, detect inherent contradictions, and calibrate confidence on difficult problems.
-  • Empirical results demonstrate synergistic improvements in both detection and reasoning.
-Crucially, we identify the phenomenon of Capability Collapse, revealing that explicit unsolvable data is essential to prevent model overconfidence.
+  • We first propose that reasoning alignment should treat objective unsolvability and subjective capability as orthogonal axes for reliable reasoning with clear decision boundaries.
+  • We introduce UnsolvableQA, a dataset built with a “Reverse Construction” pipeline that injects logical contradictions into valid rationales to create verifiable unsolvable instances.
+  • We propose UnsolvableRL, an RL paradigm with dynamic confidence thresholds and penalties for incorrect refusals, enabling models to solve feasible tasks, detect solvability, and calibrate confidence on hard inputs.
+  • Controlled ablation studies validate the necessity and synergy of our two key designs: training on unsolvable data and applying a calibrated refusal penalty.
+This combination drives a Pareto improvement, maintaining high accuracy on solvable tasks while robustly detecting unsolvable ones.
 2 Related Work
-There are two primary areas related to our work: automated dataset construction for reasoning and reinforcement learning approaches for incentivizing selective refusal.
+There are two primary concerns related to our work: automated dataset construction for reasoning and reinforcement learning approaches for incentivizing selective refusal.
+Automatic Reasoning Dataset Construction.
 As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data).
-In the realm of reasoning, recent efforts primarily focus on ensuring data solvability and verifiability through formal or programmatic means.
+In the realm of reasoning, recent efforts on automatic data construction primarily focus on ensuring data solvability and verifiability through formal or programmatic means.
 Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively.
-Concurrently, research on verifying the correctness of Chain-of-Thought (CoT) traces has also been advancing (ling2023deductive; chen2025towards).
+Concurrently, research on verifying the correctness of rationale traces has also been advancing (ling2023deductive; chen2025towards).
 In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow.
-Although earlier benchmarks like SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) attempted to include unanswerable questions, their task formats are overly simplistic (primarily extractive span selection) and fail to effectively evaluate the refusal capabilities of modern models in complex logical reasoning scenarios.
-Consequently, existing data synthesis efforts remain largely concentrated on solvable instances, leaving a significant research gap.
+Earlier, SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings.
+Consequently, for reasoning tasks, existing data-synthesis efforts focus mainly on solvable instances while largely neglecting unsolvable questions, leaving a significant research gap.
 Figure 2: The UnsolvableQA data collection pipeline.
@@
 (b) Puzzle Data Collection: We utilize domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
+Reinforcement Learning for Reasoning.
 Beyond dataset construction, Reinforcement Learning (RL) (ouyang2022traininglanguagemodelsfollow; schulman2017proximalpolicyoptimizationalgorithms; rafailov2024directpreferenceoptimizationlanguage) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness.
@@
 3 UnsolvableQA Construction
-Existing benchmarks mainly target solvable tasks, leaving models prone to hallucinations when judging feasibility.
-To address this issue, we introduce UnsolvableQA, a dataset for detecting problem inherent contradictions.
-To achieve this, we use a dual-track methodology (illustrated in Figure 2) tailored to the specific verification constraints of different domains.
+Existing benchmarks mainly target solvable tasks, so models remain prone to hallucinations when assessing feasibility.
+To address this, we introduce UnsolvableQA, a dataset for detecting intrinsic contradictions in problem statements, using a dual-track methodology (in Figure 2) aligned with domain-specific verification constraints.
 3.1 Mathematics Problem Generation via Reverse Construction
@@
 Our reverse construction pipeline operates through two sequential stages:
-Data Generation Process: First, we validate each seed problem by confirming that a strong reasoning model (DeepSeek-Reasoner) can correctly solve it and produce a complete chain-of-thought, ensuring sufficient reasoning capability.
+Data Generation Process.
+First, we validate each seed problem by confirming that a strong reasoning model (DeepSeek-Reasoner) can correctly solve it and produce a complete chain-of-thought, ensuring sufficient reasoning capability.
 Then, we employ DeepSeek-Reasoner to design logical contradictions via two complementary strategies: Constraint Contradiction, which introduces conflicting external conditions, and Axiom Contradiction, which modifies reasoning steps to violate fundamental mathematical axioms.
 The model generates a structured plan specifying the location and mechanism of contradiction injection.
-Validation Process: Based on this plan, the model then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction.
-Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-Reasoner to autonomously detect contradictions; if this fails, we explicitly provide the contradiction plan to assess genuine unsolvability.
-Only instances passing verification are retained.
+Validation Process.
+Based on this plan, LLMs then reverse-engineers a new problem statement that preserves the stylistic and complexity characteristics of the seed problem while embedding the designed contradiction.
+Finally, we implement a two-tier verification protocol: the generated problem is first presented to DeepSeek-R1 to autonomously detect contradictions; if model fails to independently identify the contradiction, we proceed to the second tier.
+In this stage, we explicitly provide the model with both the final unsolvable problem statement and the predefined contradiction plan, thereby supplementing it with the intended contradiction information beyond the problem text alone.
+The model is then prompted to reassess the problem’s solvability given this complete contextual guidance.
+Only instances that are ultimately verified as unsolvable through this two-tier procedure are retained.
 3.2 Logic Puzzle Generation
@@
 Hitori.
-We formulate Hitori as a Constraint Satisfaction Problem and, using python-constraint, generate puzzles with unique solutions.
+We formulate Hitori as a constraint satisfaction problem and, using python-constraint to generate puzzles with unique solutions.
 For unsolvable instances, we inject conflicting constraints, for example forcing a row or column to contain duplicate numbers that cannot be removed without violating adjacency rules.
@@
 Test Maze(Easy) 100 94 194
+Test Maze(Hard) 100 100 200
 Test AIME (2024–2025) 60 47 107
 Train Total 348 289 637
-Test Total 358 341 699
-Table 1: Unified vertical statistics for UnsolvableQA train/test partitions. Train Maze has no difficulty split; AIME year ranges differ. (Note: AIME year ranges: train 1983–2023; test 2024–2025.)
+Test Total 458 441 899
+Table 1: Unified vertical statistics for UnsolvableQA train/test partitions. Train Maze has no difficulty split; AIME year ranges differ.
 3.3 Dataset Statistics and Quality Control
-We construct UnsolvableQA to have approximately balanced solvable and unsolvable instances per domain, with minor residual imbalances caused by generator yields and verification constraints.
-Table 1 therefore reports aggregated statistics across the train and test partitions of UnsolvableQA.
-The test set contains 699 problems (358 solvable, 341 unsolvable), and the train set contains 637 problems (348 solvable, 289 unsolvable).
-The train partition pools all Maze instances without difficulty splits, whereas the test partition preserves Maze Easy.
-The AIME partitions cover different year ranges.
-Test prompts are diversified and paraphrased, while train prompts follow templated formats for scalable generation.
-All instances pass automated quality-control checks: logic puzzles are verified with multiple SAT/CSP/graph solvers, maze and programmatic tasks with deterministic solvers, and mathematically constructed problems with multi-pass LLM-based verification.
+We construct UnsolvableQA to maintain approximately balanced instances per domain (Table 1).
+The dataset comprises 899 test problems (458 solvable, 441 unsolvable) and 637 training problems (348 solvable, 289 unsolvable).
+While training prompts follow templated formats and pool Maze difficulties, test prompts are diversified and preserve fine-grained splits (including distinct AIME years [^1] AIME year ranges: train 1983–2023; test 2024–2025.).
+All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification.
+We also introduce two OOD testsets in Appendix A.2.
 4 UnsolvableRL Methodology
+Figure 3: UnsolvableRL framework with two orthogonal components: (a) Unsolvability Detection, which penalizes labeling solvable queries as unsolvable to preserve reasoning incentives.
+(b) Capability Calibration, which regulates refusal by comparing the target threshold ($\tau$) with rollout accuracy ($\beta$); it encourages refusal when $\tau>\beta$ and discourages it when $\tau<\beta$ via $\lambda(\tau_{i}-\beta_{i})$.
 4.1 Problem Formulation
@@
 Third, for hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
-Figure 3: UnsolvableRL Reward Framework.
-We decompose the alignment into two orthogonal components: (a) Objective Detection, which incentivizes the model to distinguish between solvable and unsolvable queries.
-Crucially, we impose a penalty (-0.5) for falsely labeling solvable problems as unsolvable to safeguard the model’s reasoning drive.
-(b) Subjective Calibration, which dynamically regulates refusal behavior.
-By comparing the target threshold ($\tau$) with the model’s rollout accuracy ($\beta$), the mechanism encourages refusal on hard tasks (where $\tau>\beta$) and suppresses it on easy ones (where $\tau<\beta$) via the dynamic reward term $\lambda(\tau_{i}-\beta_{i})$.
 4.2 Group-Relative Policy Optimization
@@
 4.3 Reward Design
+SOTA Baselines Qwen3-1.7B Qwen3-4B
+Dataset Metric Deepseek GPT-5.1 Gemini-3 Instruct + UnsolRL Instruct + UnsolRL
+Game24 S / U 98.0 / 77.5 14.0 / 50.0 94.0 / 94.0 78.0 / 23.0 84.0 / 100.0 81.0 / 17.0 94.5 / 99.0
+Mean 87.8 32.0 94.0 50.5 92.0 49.0 96.8
+HamCycle S / U 83.5 / 94.0 8.3 / 82.0 100.0 / 98.0 22.9 / 28.0 20.8 / 88.0 37.5 / 57.0 41.1 / 94.5
+Mean 88.8 45.1 99.0 25.5 54.4 47.3 67.8
+HamPath S / U 84.0 / 96.0 50.0 / 86.0 100.0 / 90.0 14.0 / 33.0 27.0 / 85.0 37.0 / 56.5 55.0 / 96.5
+Mean 90.0 68.0 95.0 23.5 56.0 46.8 75.8
+Hitori S / U 70.0 / 86.0 6.0 / 12.0 98.0 / 100.0 8.0 / 21.0 11.0 / 59.0 34.5 / 6.5 63.5 / 94.5
+Mean 78.0 9.0 99.0 14.5 35.0 20.5 79.0
+Maze (Easy) S / U 100.0 / 98.9 95.0 / 100.0 99.0 / 94.6 0.0 / 85.6 0.0 / 95.2 32.7 / 55.6 96.5 / 98.9
+Mean 99.5 97.5 96.8 42.8 47.6 44.1 97.7
+Maze (Hard) S / U 98.0 / 98.0 86.0 / 94.0 96.0 / 91.0 0.0 / 80.0 0.0 / 79.0 13.3 / 64.5 65.8 / 89.0
+Mean 98.0 90.0 93.5 40.0 39.5 38.9 77.9
+AIME 24-25 S / U 85.0 / 63.8 61.7 / 48.9 95.0 / 34.1 38.3 / 33.5 35.4 / 52.7 67.9 / 25.3 69.6 / 62.2
+Mean 74.4 55.3 64.5 35.9 44.1 46.6 65.9
+Overall S / U 88.4 / 87.7 45.9 / 67.6 97.4 / 86.0 23.0 / 43.4 25.5 / 79.8 43.4 / 40.3 69.4 / 90.7
+Mean 88.1 56.7 91.7 33.2 52.7 41.9 80.1
+Table 2: Detailed performance comparison.
+For each dataset, the first row reports Solvable Accuracy (S) / Unsolvable Detection Rate (U), and the second row reports their Mean (M).
+We bold the global best score across all models for each metric.
+Qwen3-4B + UnsolRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
 As illustrated in Figure 3, our framework explicitly distinguishes between three types of model responses: (1) Reward for Reasoning Accuracy (providing a solution), (2) Reward for Unsolvability Detection (identifying inherent contradictions), and (3) Reward for Capability Calibration (admitting capability limits).
@@
 Reward for Reasoning Accuracy ($R_{\text{acc}}$).
-This component evaluates the correctness of the model’s judgments on verifiably solvable instances.
+It evaluates the correctness of LLMs’ judgments on verifiably solvable instances.
 (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance.
@@
 (2) To prevent degenerate behavior, we impose a False Unsolvability Penalty ($\rho=-0.5$): if the model incorrectly declares a solvable problem unsolvable, it receives a negative reward.
-We explicitly motivate this penalty value via a decision-theoretic analysis (formal proof in Appendix A.1).
+As derived in our decision-theoretic analysis in Appendix A.3, this penalty is mathematically critical to prevent the model from collapsing into a local optimum of universal rejection.
+It effectively expands the decision threshold, encouraging the model to attempt reasoning even under moderate uncertainty.
+(3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.
+Reward for Difficulty Calibration ($R_{\text{cal}}$).
+When the model outputs the specific refusal token (e.g., beyond my capabilities), we regulate refusal via a target accuracy threshold $\tau\in[0,1]$.
+During each RL rollout, we compute the empirical accuracy $\beta$ of the current group:
+$\beta=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[y_{i}\text{ is correct}].$ (3)
+The calibration reward is then defined as:
+$R_{\text{cal}}(y)=\lambda(\tau-\beta)\,\mathbf{1}[y\text{ is a refusal}].$ (4)
+The term $(\tau-\beta)$ quantifies the gap between desired and actual reliability.
+This yields a dynamic incentive: when the model’s reasoning capability is insufficient ($\beta<\tau$), the reward becomes positive, promoting refusal; conversely, when performance is strong ($\beta>\tau$), it turns negative, thus discouraging refusal on tasks within its capability.
+To enforce robust calibration, we implement a progressive schedule for $\tau$, increasing it monotonically throughout training to approach $1$.
+This curriculum balances exploration (low $\tau$) and prudent strictness (high $\tau$).
+We provide a formal proof of this Capability-Refusal Trade-off in Appendix A.4, mathematically demonstrating why a static threshold leads to reward collapse and failure to refuse.
+The final per-trajectory reward is the sum of these components:
+$R=R_{\text{acc}}+R_{\text{detect}}+R_{\text{cal}}.$ (5)
+4.4 Answer Extraction and Verification
+Solvable tasks employ deterministic verifiers, including graph/constraint solvers and reference matchers.
+Unsolvable tasks are validated by the canonical unsolvable marker.
+Finally, responses containing “beyond my capabilities” trigger the dynamic calibration reward.
+5 Experiments
+Figure 4: Detailed Ablation Metrics.
+Breakdown of Solvable Accuracy (S, Red), Unsolvable Detection Rate (U, Blue), and Mean Performance (M, Green) across different configurations.
+The w/o UnsData w/ P setting shows a catastrophic drop in U for the 1.7B model (4.2%), indicating capability collapse.
+Notably, w/ UnsData w/o P achieves the highest Mean (M) score by maximizing U.
+The full method (w/ UnsData w/ P) demonstrates a valuable trade-off, boosting S while maintaining a high U, optimizing for practical scenarios.
+Figure 5: Mechanism Analysis: Trade-off between $S$ and $U$.
+Visualizing the interaction between Data and Penalty.
+The Purple Dashed Arrow highlights the comprehensive enhancement: compared to the naive baseline (Red Square), our full method (Blue Circle) achieves significant improvements in both Solvable Accuracy ($S$) and Unsolvable Detection ($U$).
+Specific dynamics include: (left) Capability Collapse: Applying the penalty without data causes the 1.7B model to collapse (Red Arrow).
+(right) Strategic Trade-off: With adequate data, the penalty acts as a regularizer (Blue Arrow), trading a small amount of $U$ for a gain in $S$ to optimize for real-world applications.
+We trained and evaluated Qwen3-1.7B and Qwen3-4B (yang2025qwen3technicalreport) in think mode on UnsolvableQA.
+For evaluation, we report average@8 on the AIME partitions and average@4 on other datasets.
+We standardize training steps: ablation variants are trained for 120 steps, while the final models are trained for 320 steps.
+5.1 Main Results: Bridging the Gap with SOTA
+Table 2 compares our Qwen3-UnsolRL models against both proprietary SOTA models and base baselines.
+The complete experimental results are presented in Appendix 4.
+High-reasoning LLMs (e.g., Deepseek-V3.2-R1) inherently perform well, whereas base models struggle significantly (e.g., Qwen3-1.7B Instruct achieves only 33.2% mean score).
+Our method bridges this gap: Qwen3-4B + UnsolRL achieves a global mean score of 80.1%, significantly outperforming GPT-5.1-low (56.7%) and approaching the performance of Gemini-3 (91.7%).
+Notably, our 4B model achieves the highest global Unsolvable Detection rate (90.7%), surpassing even the much larger Deepseek and Gemini models.
+This confirms that equipping models with the ability to identify unsolvability ($\mathcal{U}$) strengthens the rigor required to solve feasible problems ($\mathcal{S}$).
+Figure 6: Performance Comparison and Calibration Analysis.
+(Left/Middle) Comparison with TruthRL.
+While TruthRL achieves higher rejection rates (Blue bars), it suffers from over-refusal, significantly degrading Accuracy (Green bars).
+(Right) Comparison of threshold strategies.
+The adaptive mechanism (w/ P Final) maintains a healthier rejection rate (6.87%) compared to the Fixed-$\tau$ baseline (4.36%) while preserving high accuracy.
+5.2 Ablation Study: Disentangling Data and Penalty
+To understand the contributions of data and reward mechanisms, we conducted a comprehensive ablation study comparing four configurations.
+The detailed metrics are shown in Figure 4, and the underlying dynamics are visualized in Figure 5.
+Explicit Unsolvable Data is a Prerequisite for Detection.
+As shown in Figure 4, the naive baseline trained without unsolvable data (w/o UnsData) lacks the fundamental capability to recognize unsolvability.
+For Qwen3-4B, detection is weak ($U=64.1\%$).
+Introducing explicit unsolvable data provides the necessary supervision, boosting detection significantly ($U\to 89.3\%$).
+This confirms that instruction following alone is insufficient; explicit data is the foundational requirement for boundary learning.
+Applying Penalty without Data Triggers Capability Collapse.
+Figure 5(left) illustrates a critical failure mode caused by objective mismatch.
+When the penalty is enforced without providing ground-truth unsolvable examples, the model, lacking positive examples of valid refusal, minimizes the penalty risk by defaulting to answering all queries.
+This behavior effectively collapses its refusal mechanism, as seen in the 1.7B model’s detection rate plummeting from 70.8% to 4.2%.
+Applying Penalty with Data Facilitates a Strategic Trade-off.
+When grounded with adequate data, the penalty shifts from being a suppressor to a regularizer.
+For Qwen3-4B, adding the penalty (w/ UnsData w/ P) improves Solvable Accuracy ($S$) from 60.6% to 62.6%, at the cost of a slight drop in detection ($U:89.3\%\to 84.8\%$).
+This trade-off is strategic: it sacrifices marginal detection performance to enforce stricter reasoning rigor on feasible problems, optimizing the model for real-world utility.
+Overall Gain: A Pareto Improvement over the Baseline.
+Viewing the trajectory globally (the purple dashed arrow in Figure 5b), our full method represents a comprehensive upgrade over the naive baseline (w/o UnsData w/o P).
+We achieve significant gains in both dimensions for Qwen3-4B ($S:57.7\%\to 62.6\%$, $U:64.1\%\to 84.8\%$).
+This confirms that UnsolvableRL does not merely shift the decision boundary, but fundamentally enhances both reliability and reasoning capability.
+5.3 Comparative Analysis: TruthRL and Calibration
+We further benchmark our method against TruthRL(wei2025truthrl) and different thresholding strategies. The results are summarized in Figure 6.
+Superior Balance Compared to TruthRL.
+As shown in the first two panels of Figure 6, TruthRL tends to be overly conservative, leading to over-refusal.
+For Qwen3-1.7B, while TruthRL achieves a high Rejection Rate (18.58%), its Accuracy drops to 40.4%, which is inferior to even the baseline.
+In contrast, our method achieves a significantly higher Accuracy (50.3%) with a moderate and precise Rejection Rate (7.96%).
+This indicates that our False Unsolvability Penalty provides a softer, more distinctive signal than TruthRL’s hard refusal reward, preserving the model’s core reasoning capabilities.
+Dynamic Thresholding Prevents Overconfidence.
+The rightmost panel of Figure 6 highlights the importance of our dynamic calibration.
+The Fixed-$\tau$ baseline yields a low rejection rate (4.36%), suggesting overconfidence.
+By dynamically adjusting $\tau$ based on the moving average of solvable confidence, our method (w/ P Final) maintains a robust rejection rate (6.87%) without sacrificing accuracy (80.1%), ensuring the model remains prudent even as it becomes stronger.
+6 Conclusion
+In this paper, we address the challenge of LLM reliability by formally decoupling objective unsolvability from subjective capability.
+We introduce UnsolvableQA for rigorous evaluation and UnsolvableRL, an alignment framework featuring dynamic calibration.
+Empirically, our method achieves a Pareto improvement over baselines: it boosts solvable accuracy (Overall $S$) from 43.4% to 69.4% while maintaining robust unsolvability detection ($U>90\%$ on Qwen3-4B).
+Crucially, our analysis uncovers the dual nature of alignment penalties: without grounding data, they trigger Capability Collapse; yet with adequate data, they facilitate a Strategic Trade-off, acting as a Regularizer for Rigor that prevents hallucination without the over-refusal pitfalls observed in prior methods like TruthRL.
+We hope this work establishes a new paradigm for building models that rigorously distinguish between inherent unsolvability and problems that simply exceed their current capabilities.
+Limitations
+While the UnsolvableQA dataset is relatively comprehensive within the selected domains (logic puzzles, mathematics, mazes), it may not fully encompass the various forms of inherently unsolvable problems encountered in real-world applications.
+Future work could expand the dataset to include more diverse and ambiguous types of unsolvable problems.
+Appendix A Appendix
+A.1 Comprehensive experimental setup and results
+Game24 HamCycle HamPath Hitori Maze(E) Maze(H) AIME24-25 Overall
+Checkpoint S U M S U M S U M S U M S U M S U M S U M $\mathrm{S}$ $\mathrm{U}$ $\mathrm{M}$
+Deepseek-V3.2-R 98.0 77.5 87.8 83.5 94.0 88.8 84.0 96.0 90.0 70.0 86.0 78.0 100.0 98.9 99.5 98.0 98.0 98.0 85.0 63.8 74.4 88.4 87.7 88.1
+GPT-5.1-Low 14.0 50.0 32.0 8.3 82.0 45.1 50.0 86.0 68.0 6.0 12.0 9.0 95.0 100.0 97.5 86.0 94.0 90.0 61.7 48.9 55.3 45.9 67.6 56.7
+Gemini-3.0-Pro 94.0 94.0 94.0 100.0 98.0 99.0 100.0 90.0 95.0 98.0 100.0 99.0 99.0 94.6 96.8 96.0 91.0 93.5 95.0 34.1 64.5 97.4 86.0 91.7
+Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 13.3 64.5 38.9 67.9 25.3 46.6 43.4 40.3 41.9
++ TruthRL 91.0 91.0 91.0 32.3 86.0 59.2 49.0 90.0 69.5 38.0 15.0 26.5 90.0 77.1 83.6 31.0 56.5 43.8 67.5 25.5 46.5 57.0 63.0 60.0
++ w/o UnsData w/o P 83.0 92.0 87.5 36.5 81.0 58.8 47.0 84.0 65.5 46.0 20.0 33.0 88.5 76.6 82.6 34.5 60.5 47.5 68.3 34.6 51.5 57.7 64.1 60.9
++ w/o UnsData w/ P 90.5 12.5 51.5 39.6 67.5 53.5 49.5 64.5 57.0 56.5 4.0 30.3 94.8 83.2 89.0 56.3 72.0 64.1 69.0 18.4 43.7 65.2 46.0 55.6
++ w/ UnsData w/o P 84.0 95.0 89.5 39.6 96.0 67.8 53.0 100.0 76.5 45.0 87.0 66.0 95.5 95.2 95.4 40.0 89.5 64.8 67.1 62.2 64.7 60.6 89.3 75.0
++ w/ UnsData w/ P 88.0 97.0 92.5 35.9 95.5 65.7 51.5 99.5 75.5 48.5 71.0 59.8 94.0 95.5 94.8 53.3 79.8 66.6 66.9 55.3 61.1 62.6 84.8 73.7
++ Fixed-$\tau$ Final 87.5 99.0 93.3 39.1 100.0 69.6 67.0 99.5 83.3 51.0 98.0 74.5 96.3 100.0 98.2 52.5 95.8 74.2 69.0 63.6 66.3 66.1 93.7 79.9
++ w/ UnsData w/ P Final 94.5 99.0 96.8 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.5 98.9 97.7 65.8 89.0 77.9 69.6 62.2 65.9 69.4 90.7 80.1
+Qwen3-1.7B-Instruct 78.0 23.0 50.5 22.9 28.0 25.5 14.0 33.0 23.5 8.0 21.0 14.5 0.0 85.6 42.8 0.0 80.0 40.0 38.3 33.5 35.9 23.0 43.4 33.2
++ TruthRL 76.0 80.0 78.0 19.8 49.0 34.4 17.0 58.0 37.5 4.0 25.0 14.5 1.5 82.4 42.0 0.0 73.5 36.8 42.1 37.2 39.7 22.9 57.9 40.4
++ w/o UnsData w/o P 77.0 88.0 82.5 25.0 60.0 42.5 15.0 83.0 49.0 5.0 53.0 29.0 0.5 93.6 47.1 0.0 79.5 39.8 32.9 38.3 35.6 22.2 70.8 46.5
++ w/o UnsData w/ P 85.0 2.0 43.5 22.4 10.5 16.5 25.5 14.0 19.8 12.0 2.5 7.3 0.0 0.0 0.0 0.0 0.0 0.0 40.8 0.5 20.7 26.5 4.2 15.4
++ w/ UnsData w/o P 70.0 97.0 83.5 22.9 94.0 58.5 22.0 97.0 59.5 4.0 79.0 41.5 0.5 95.2 47.9 0.0 86.0 43.0 31.7 56.4 44.0 21.6 86.4 54.0
++ w/ UnsData w/ P 82.0 96.5 89.3 18.2 85.0 51.6 20.5 89.5 55.0 8.0 62.0 35.0 0.0 90.7 45.4 0.0 74.3 37.2 40.8 36.4 38.6 24.2 76.3 50.3
++ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2
+Table 3: Detailed performance breakdown (S/U/M) across all seven datasets.
+This table lists the full experimental results for all baselines and Qwen3 variants (including ablation checkpoints and final converged models).
+S: Solvable Accuracy; U: Unsolvable Detection Rate; M: Mean Score.
+Game24 HamCycle HamPath Hitori Maze(E) Maze(H) AIME24-25 Overall
+Checkpoint Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R
+Deepseek-V3.2-R 87.8 0.00 87.8 88.8 8.30 97.1 90.0 6.00 96.0 78.0 9.00 87.0 99.5 0.00 99.5 98.0 0.00 98.0 74.4 0.80 75.2 88.1 3.44 91.5
+GPT-5.1-Low 32.0 1.00 33.0 45.1 2.08 47.2 68.0 8.00 76.0 9.0 1.00 10.0 97.5 0.00 97.5 90.0 0.00 90.0 55.3 7.90 63.2 56.7 2.85 59.6
+Gemini-3.0-Pro 94.0 1.00 95.0 99.0 0.00 99.0 95.0 0.00 95.0 99.0 0.00 99.0 96.8 0.00 96.8 93.5 0.00 93.5 64.5 0.00 64.5 91.7 0.14 91.8
+Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 38.9 0.25 39.2 46.6 0.00 46.6 41.9 5.86 47.7
++ TruthRL 91.0 0.50 91.5 59.2 34.15 93.4 69.5 16.00 85.5 26.5 18.50 45.0 83.6 2.50 86.1 43.8 11.00 54.8 46.5 0.00 46.5 60.0 11.81 71.8
++ w/o UnsData w/o P 87.5 0.00 87.5 58.8 28.85 87.7 65.5 15.50 81.0 33.0 11.00 44.0 82.6 0.00 82.6 47.5 0.00 47.5 51.5 0.00 51.5 60.9 7.91 68.8
++ w/o UnsData w/ P 51.5 0.00 51.5 53.5 21.09 74.6 57.0 7.00 64.0 30.3 2.25 32.6 89.0 0.00 89.0 64.1 0.13 64.2 43.7 0.00 43.7 55.6 4.35 60.0
++ w/ UnsData w/o P 89.5 1.50 91.0 67.8 19.00 86.8 76.5 10.00 86.5 66.0 13.00 79.0 95.4 1.00 96.4 64.8 4.00 68.8 64.7 0.42 65.1 75.0 7.00 82.0
++ w/ UnsData w/ P 92.5 1.00 93.5 65.7 19.75 85.5 75.5 15.50 91.0 59.8 7.50 67.3 94.8 0.00 94.8 66.6 1.00 67.6 61.1 0.00 61.1 73.7 6.39 80.1
++ Fixed-$\tau$ Final 93.3 0.00 93.3 69.6 18.49 88.1 83.3 12.00 95.3 74.5 0.00 74.5 98.2 0.00 98.2 74.2 0.00 74.2 66.3 0.00 66.3 79.9 4.36 84.3
++ w/ UnsData w/ P Final 96.8 0.00 96.8 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.7 0.00 97.7 77.1 0.00 77.1 65.9 0.00 65.9 80.1 6.87 87.0
+Qwen3-1.7B-Instruct 50.5 0.00 50.5 25.5 8.85 34.3 23.5 12.00 35.5 14.5 0.50 15.0 42.8 0.00 42.8 40.0 0.00 40.0 35.9 0.00 35.9 33.2 3.05 36.3
++ TruthRL 78.0 2.00 80.0 34.4 29.58 64.0 37.5 39.50 77.0 14.5 16.50 31.0 42.0 12.78 54.8 36.8 24.00 60.8 39.7 5.68 45.4 40.4 18.58 59.0
++ w/o UnsData w/o P 82.5 0.50 83.0 42.5 29.17 71.7 49.0 30.00 79.0 29.0 6.50 35.5 47.1 0.25 47.3 39.8 3.50 43.3 35.6 4.17 39.8 46.5 10.58 57.1
++ w/o UnsData w/ P 43.5 0.25 43.8 16.5 19.01 35.5 19.8 8.00 27.8 7.3 5.25 12.5 0.0 0.00 0.0 0.0 0.00 0.0 20.7 0.00 20.7 15.4 4.64 20.1
++ w/ UnsData w/o P 83.5 1.00 84.5 58.5 22.88 81.4 59.5 19.50 79.0 41.5 2.50 44.0 47.9 1.50 49.4 43.0 0.75 43.8 44.0 2.52 46.5 54.0 7.24 61.2
++ w/ UnsData w/ P 89.3 0.00 89.3 51.6 27.77 79.4 55.0 19.50 74.5 35.0 2.50 37.5 45.4 0.00 45.4 37.2 4.50 41.7 38.6 1.46 40.1 50.3 7.96 58.3
++ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2
+Table 4: Alternative metric analysis showing Correctness (Corr), Rejection Rate (Rej), and their sum (C+R).
+This breakdown distinguishes between general accuracy and refusal tendency, providing a diagnostic view for over-refusal and capability collapse.
+In this appendix, we provide the full, domain-level experimental results to supplement the aggregated metrics reported in the main text.
+Table 3 reports the performance breakdown across all seven datasets, including Game24, HamCycle, HamPath, Hitori, Maze-Easy, Maze-Hard, and AIME 24-25.
+We report three primary metrics to evaluate performance.
+First, Solvable Accuracy (S) measures the percentage of solvable problems correctly answered.
+Second, Unsolvable Detection Rate (U) indicates the percentage of unsolvable problems correctly identified as unsolvable.
+Finally, Mean (M) represents the arithmetic mean of S and U, reflecting the overall reliability of the model across different domains.
+Complementing the standard metrics, Table 4 provides a diagnostic view focusing on the model’s refusal behavior.
+Correctness (Corr) measures the accuracy of the model’s responses on attempted questions, independent of the refusal decision.
+Rejection Rate (Rej) tracks the proportion of total queries, including both solvable and unsolvable instances, that the model refused to answer.
+The sum of these two metrics (C+R) serves as a diagnostic tool to analyze trade-offs between reasoning capability and refusal tendency, helping to identify issues such as over-refusal or capability collapse.
+Both tables include results for a wide range of configurations.
+We compare against strong proprietary models like Deepseek-V3.2-R, GPT-5.1-Low, and Gemini-3.0-Pro, as well as the base Qwen3-Instruct models, which serve as reference points.
+To isolate the effects of our proposed components, we evaluate ablation variants trained for 120 steps.
+These include models trained exclusively on solvable data without exposure to ground-truth unsolvable examples (denoted as w/o UnsData) and models trained without the False Unsolvability Penalty (denoted as w/o P), representing standard RLHF without the penalty term for incorrect refusals.
+Finally, we present results for the fully converged models trained for 320 steps, including a variant using a static confidence threshold (Fixed-tau Final) and our fully converged model trained with the complete UnsolvableRL method (w/ UnsData w/ P Final).
+A.2 Out-of-Distribution (OOD) Generalization
+To evaluate the generalization capabilities of UnsolRL, we utilized two Out-of-Distribution (OOD) benchmarks: 8-Puzzle ($N_{sol}=87,N_{unsol}=100$) and Zebra Logic ($N_{sol}=100,N_{unsol}=100$).
+These tasks differ significantly from the training distribution, serving as a rigorous test for reasoning robustness.
+Table 5 details the performance mean@4 across Solvable (S), Unsolvable (U), and Mean (M) metrics.
+SOTA Model Performance.
+We first established a baseline using top-tier proprietary models.
+Deepseek-V3.2-R achieves the highest overall performance (M=76.2%), demonstrating exceptional dominance in logic tasks with a mean score of 92.7% on Zebra Logic.
+In contrast, Gemini-3.0-Pro excels in procedural planning, securing the top spot on the 8-Puzzle benchmark (M=70.3%).
+GPT-5.1-Low shows balanced capabilities, particularly in detecting unsolvable instances across both domains (Overall U=94.5%).
+These results highlight that while leading models perform well, handling both logical constraints and path-finding robustly remains a non-trivial challenge.
+Qwen3-4B Results.
+The baseline Qwen3-4B model struggles significantly with OOD detection, failing completely to identify unsolvable 8-Puzzle instances (0.0%).
+UnsolRL-Final substantially enhances robustness: it raises the U score to 15.8% on the 8-Puzzle and achieves a remarkable improvement on Zebra Logic, boosting U from 69.0% to 87.0%.
+Consequently, the Overall Mean (M) increases by over 50% relative to the baseline (from 23.2% to 34.9%), demonstrating that our method successfully transfers verification capabilities to unseen tasks.
+Qwen3-1.7B Results.
+Despite its limited capacity, the 1.7B model exhibits a similar positive trend in anomaly detection.
+On Zebra Logic, UnsolRL improves the unsolvable detection rate (U) from 36.0% to 47.3%.
+Although the complex 8-Puzzle remains highly challenging for this smaller architecture (Overall M for 8-Puzzle improves marginally to 1.1%), the improvement in the Overall Mean across tasks (from 11.5% to 13.8%) confirms that UnsolRL effectively mitigates hallucinations on unsolvable queries, even when reasoning resources are constrained.
+8-Puzzle Zebra Logic Overall
+Model S U M S U M S U M
+Deepseek-V3.2-R 40.2 79.0 59.6 94.5 91.0 92.7 67.3 85.0 76.2
+Gemini-3.0-Pro 43.6 97.0 70.3 65.0 74.0 69.5 54.3 85.5 69.9
+GPT-5.1-Low 28.9 98.0 63.5 48.0 91.0 69.5 38.5 94.5 71.5
+Qwen3-4B-Instruct 21.6 0.0 10.8 2.0 69.0 35.5 11.8 34.5 23.2
++ UnsolRL-Final 26.4 15.8 20.1 8.8 87.0 47.9 17.4 52.4 34.9
+Qwen3-1.7B-Instruct 0.8 0.0 0.4 9.0 36.0 22.5 4.9 18.0 11.5
++ UnsolRL-Final 1.2 1.0 1.1 5.3 47.3 26.3 3.3 24.2 13.8
+Table 5: OOD performance. S: Solvable, U: Unsolvable, M: Mean accuracy (%).
+A.3 Theoretical Justification for False Unsolvability Penalty
+We explicitly motivate the penalty value $\rho=-0.5$ via a decision-theoretic analysis.
 Consider a model with an initial weak reasoning capability $\epsilon$ on solvable tasks (e.g., $\epsilon\approx 0.1$) and a posterior belief $p$ that a given problem is unsolvable.
@@
 This “Threshold Expansion” effectively encourages the model to attempt reasoning even under moderate uncertainty, preventing the collapse of reasoning capabilities.
-(3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.
-Game24 HamCycle HamPath Hitori Maze(E) AIME24-25 Overall
-S U M S U M S U M S U M S U M S U M $\mathrm{S}$ $\mathrm{U}$ $\mathrm{M}$
-Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 67.9 25.3 46.6 48.4 36.3 42.4
-+ Solvable-Only 85.5 11.5 48.5 41.1 65.0 53.1 52.0 66.5 59.3 53.5 5.0 29.3 95.0 83.0 89.0 68.3 18.9 43.6 65.9 41.7 53.8
-+ UnsolvableRL-Mid 84.5 98.5 91.5 36.5 95.5 66.0 44.0 98.0 71.0 47.0 72.5 59.8 95.7 94.9 95.3 67.3 52.7 60.0 62.5 85.4 73.9
-+ Fixed-$\tau$ 85.5 99.0 92.3 43.8 99.5 71.6 59.5 99.0 79.3 48.5 98.0 73.3 96.6 99.5 98.0 69.6 61.4 65.5 67.3 92.7 80.0
-+ UnsolvableRL-Final 92.0 99.0 95.5 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.0 98.9 97.5 69.2 85.0 77.1 69.5 90.9 80.3
-Qwen3-1.7B-Instruct 63.0 21.0 42.0 22.9 22.5 22.7 13.0 26.5 19.8 6.0 27.0 16.5 3.6 91.0 47.3 38.3 37.5 37.9 24.5 37.6 31.0
-+ Solvable-Only 78.5 1.0 39.8 23.4 2.0 12.7 13.5 2.5 8.0 4.5 1.5 3.0 5.7 0.0 2.9 41.7 2.1 21.9 27.9 1.5 14.7
-+ UnsolvableRL-Mid 70.5 90.5 80.5 19.3 85.5 52.4 17.0 91.5 54.3 11.0 56.0 33.5 4.1 92.0 48.1 39.2 39.6 39.4 26.9 75.9 51.4
-+ UnsolvableRL-Final 79.0 99.5 89.3 21.9 87.5 54.7 17.0 87.5 52.3 6.0 56.5 31.3 8.6 96.8 52.7 37.3 48.4 42.8 28.3 79.4 53.9
-Table 2: Per-domain performance for Qwen3 checkpoints.
-Values are percentages.
-For each test domain we report solvable accuracy (S), unsolvable rejection rate (U), and their mean (M = (S+U)/2).
-Overall metrics are unweighted means across all 6 domains.
-Model names indicate the training configuration and stage: UnsolvableRL is our main method trained on paired solvable/unsolvable data; Solvable-Only is an ablation trained exclusively on solvable instances (Step 60 for 1.7B, Step 120 for 4B); Fixed-$\tau$ is an ablation using a fixed calibration threshold.
-Suffixes Mid and Final represent intermediate (Step 60 for 1.7B, Step 120 for 4B) and extended (Step 360 for 1.7B, Step 320 for 4B) trained models, respectively.
-Reward for Subjective Calibration ($R_{\text{cal}}$).
-When the model outputs the specific refusal token (e.g., beyond my capabilities), we do not apply a static score.
-Instead, we regulate refusal via a target accuracy threshold $\tau\in[0,1]$.
-During each RL rollout, we compute the empirical accuracy $\beta$ of the current group as:
-$\beta=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}[y_{i}\text{ is correct}],$ (3)
-where correctness includes both solving solvable problems and identifying unsolvable ones. The calibration reward is then defined as:
-$R_{\text{cal}}(y)=\lambda(\tau-\beta)\,\mathbf{1}[y\text{ is a refusal}].$ (4)
-Here, $(\tau-\beta)$ quantifies the reliability deficit—the gap between the target and the model’s actual performance.
-This creates a dynamic incentive mechanism as shown in Figure 3(b): when the model’s reasoning capability on the current batch is poor ($\beta<\tau$), the term is positive, encouraging refusal; conversely, when performance is high ($\beta>\tau$), the reward becomes negative, suppressing refusal on easy questions.
-To enforce robust calibration, we implement a progressive schedule for $\tau$, increasing it steadily throughout training to approach $1$.
-This curriculum balances exploration and caution: in early stages, a lower $\tau$ provides a lenient margin to encourage the exploration of reasoning paths; in later stages, as $\tau\to 1$, it enforces a strict standard, compelling the model to answer prudently.
-We mathematically justify the necessity of this dynamic schedule by analyzing the general condition for refusal.
+A.4 Theoretical Analysis of Capability-Refusal Dynamics
+In this section, we derive the mathematical relationship between the model’s reasoning capability and its tendency to refuse, explaining why a dynamic threshold is essential.
 The model should choose to refuse only if the reward for refusal exceeds the expected reward of answering:
@@
 Note that the expected reward for answering is always non-negative ($\mathbb{E}[R_{\text{ans}}]\geq 0$).
-If we were to use a fixed threshold (e.g., $\tau=0.5$), as the model’s reasoning capability improves, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$).
-Once this occurs, the refusal reward becomes negative ($R_{\text{cal}}<0$).
-In this regime, the inequality $R_{\text{cal}}>\mathbb{E}[R_{\text{ans}}]$ becomes impossible to satisfy, as a negative reward cannot exceed a non-negative one.
-Consequently, for well-performed batches, the model is incentivized to suppress refusal entirely to avoid penalties.
+Failure of Fixed Threshold.
+If we were to use a fixed threshold (e.g., $\tau=0.5$), as the model’s reasoning capability improves during training, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$).
+Once this occurs, the refusal reward becomes negative ($R_{\text{cal}}=\lambda(\tau-\beta)<0$).
+Substituting this into Inequality 6, the condition becomes $\text{Negative Value}>\mathbb{E}[R_{\text{ans}}]$.
+Since $\mathbb{E}[R_{\text{ans}}]\geq 0$, this inequality is impossible to satisfy.
+Consequently, for well-performed batches, the model is incentivized to suppress refusal entirely to avoid penalties, effectively forcing the model to hallucinate answers on difficult problems rather than refusing.
 Our progressive schedule prevents this collapse by ensuring $\tau$ grows dynamically to maintain $\tau\gtrsim\beta$, keeping the refusal incentive positive for uncertain queries.
-The final per-trajectory reward is the sum of these components:
-$R=R_{\text{acc}}+R_{\text{detect}}+R_{\text{cal}}.$ (6)
-Game24 HamCycle HamPath Hitori Maze(E) AIME24-25 Overall
-Checkpoint Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R
-Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 46.6 0.00 46.6 42.4 6.79 49.2
-+ Solvable-Only 48.5 0.00 48.5 53.1 20.57 73.7 59.3 7.50 66.8 29.3 1.50 30.8 89.0 0.00 89.0 43.6 0.00 43.6 53.8 4.93 58.7
-+ UnsolvableRL-Mid 91.5 0.00 91.5 66.0 19.01 85.0 71.0 13.00 84.0 59.8 1.50 61.3 95.3 0.00 95.3 60.0 0.00 60.0 73.9 5.59 79.5
-+ Fixed-$\tau$ 92.3 0.00 92.3 71.6 17.19 88.8 79.3 10.00 89.3 73.3 1.00 74.3 98.0 0.00 98.0 65.5 0.00 65.5 80.0 4.70 84.7
-+ UnsolvableRL-Final 95.5 0.00 95.5 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.5 0.00 97.5 65.9 0.00 65.9 80.3 8.01 88.3
-Qwen3-1.7B-Instruct 42.0 0.00 42.0 22.7 9.38 32.1 19.8 9.25 29.0 16.5 0.50 17.0 47.3 0.00 47.3 37.9 0.21 38.1 31.0 3.22 34.3
-+ Solvable-Only 39.8 0.00 39.8 12.7 10.94 23.6 8.0 7.00 15.0 3.0 0.25 3.3 2.9 0.00 2.9 21.9 0.10 22.0 14.7 3.05 17.8
-+ UnsolvableRL-Mid 80.5 0.00 80.5 52.4 21.35 73.8 54.3 13.00 67.3 33.5 2.00 35.5 48.1 0.00 48.1 39.4 2.08 41.5 51.4 6.41 57.8
-+ UnsolvableRL-Final 89.3 0.25 89.6 54.7 23.44 78.1 52.3 18.00 70.3 31.3 1.75 33.0 52.7 0.00 52.7 42.8 1.88 44.7 53.9 7.55 61.4
-Table 3: Aggregated domain-level metrics for Qwen3 checkpoints.
-Each domain reports: Correctness (mean of solvable and unsolvable correct rates), Rejection (mean rejection rate), and Combined = Correctness + Rejection.
-Overall columns are unweighted means across domains.
-4.4 Answer Extraction and Verification
-For solvable tasks, deterministic verifiers (graph and constraint solvers, arithmetic evaluators, and reference-answer matchers) are applied.
-For unsolvable tasks, verification instead checks for the canonical unsolvable marker.
-For logic puzzles, domain-specific verifiers check the extracted answer against problem constraints and ground-truth solutions.
-For AIME and other mathematics problems, “math_verify” is used for symbolic equivalence checking.
-Answers containing “beyond my capabilities” are treated as calibration and handled by the dynamic reward logic.
-5 Experiments
-We evaluated Qwen3-1.7B and Qwen3-4B (yang2025qwen3technicalreport) with think mode after training and testing on UnsolvableQA.
-For evaluation, we report average@8 on the AIME solvable and unsolvable partitions to account for higher variance, and average@4 on all other datasets.
-Figure 4: Evolution of Unsolvable Accuracy during training.
-The curves represent the average detection accuracy across unsolvable instances from the evaluated domains.
-Models trained with UnsolvableQA (solid lines) steadily improve their ability to identify contradictions.
-In contrast, the “No Data” ablation (dashed lines) leads to stagnation or, specifically for the 1.7B model, a catastrophic collapse in unsolvability detection.
-UnsolvableRL substantially improves both accuracy and calibration.
-As shown in Table 3, fine-tuning on a mixture of solvable and unsolvable data yields a consistent increase in overall accuracy across all test sets for both models.
-Aggregated metrics indicate that Qwen3-4B’s combined performance score (C+R) increases from 49.2% to 88.3% under the UnsolvableRL-Final configuration.
-Crucially, UnsolvableRL significantly improves unsolvable detection capabilities: the mean rejection rate on unsolvable instances ($U_{\mathrm{mean}}$) surged from 36.3% to 90.9% for Qwen3-4B, and from 37.6% to 79.4% for Qwen3-1.7B.
-In complex reasoning domains such as Game24 and Hamiltonian Path, the model attains rejection rates above 94% on unsolvable instances while maintaining, and in some cases improving, performance on solvable ones; for example, solvable accuracy on Game24 improves from 81.0% to 92.0%.
-Together, these results demonstrate that our training methodology enables the models to reliably distinguish between solvable and unsolvable problems and to exhibit the desired behavior of either answering correctly or abstaining.
-Training on solvable-only data is fundamentally insufficient for learning robust calibration.
-We further conducted an ablation study by training models on a variant of our data with all unsolvable instances removed (Solvable-Only).
-We compare the Solvable-Only model with the full-training (UnsolvableRL) model at the same RL steps.
-As shown in Table 2, the Qwen3-1.7B-Solvable-Only model’s mean unsolvable rejection rate collapses to 1.5%, compared to 75.9% for the UnsolvableRL model.
-This catastrophic failure underscores the necessity of explicitly including unsolvable examples during training to prevent over-confidence.
-Dynamic thresholds achieve a more favorable trade-off between performance and calibration.
-To assess the effectiveness of the Reward for Subjective Calibration, we compare our main model with a Fixed-$\tau$ variant that uses a static low threshold.
-As shown in Table 3, the dynamic mechanism substantially increases refusal under uncertainty, raising the mean rejection rate on solvable instances from 4.70% (Fixed-$\tau$) to 8.01%.
-This behavior arises from the interaction between model capability and reward.
-As training progresses and the model’s reasoning capability improves, the empirical accuracy $\beta$ increases.
-Under Fixed-$\tau$, once performance exceeds the static threshold ($\beta>\tau$), the calibration reward term $\lambda(\tau-\beta)$ becomes negative, penalizing refusal and effectively forcing the model to hallucinate answers on difficult problems.
-In contrast, our dynamic mechanism adjusts $\tau$ based on $\beta$, maintaining a positive incentive for refusal only when necessary.
-We provide a formal proof of this Capability-Refusal Trade-off in Appendix A.2.
-Consequently, the dynamic approach achieves a better balance, maintaining strong accuracy (Overall $S$ of 69.5% vs.
-67.3%) alongside safer behavior.
-Training without unsolvable instances leads to the collapse of unsolvability detection.
-To further assess the role of negative supervision, we analyze the training dynamics in Figure 4.
-We track the Unsolvable Accuracy, defined as the accuracy of detecting unsolvable instances averaged across our evaluation domains.
-Models trained with our paired data (solid lines) steadily improve at detecting inherent contradictions, whereas ablation models without unsolvable samples (dashed lines) fail to generalize.
-This gap is most pronounced in the 1.7B model, which exhibits a sharp capability collapse, with its accuracy on unsolvable problems approaching zero.
-We attribute this to Gradient Interference in smaller models, where optimization for answering solvable questions suppresses the features required for detection (see Appendix A.3 for a theoretical derivation).
-These results demonstrate that explicit exposure to unsolvable problems is crucial for preventing overconfidence.
-6 Conclusion
-In this paper, we enhance the reliability of large language models (LLMs) by introducing two novel perspectives: detecting inherently unsolvable problems and calibrating model behavior on problems beyond their capabilities.
-To this end, we construct UnsolvableQA using a novel Reverse Construction methodology and align models with UnsolvableRL, which incorporates a dynamic refusal mechanism.
-Our approach enables models to effectively refuse both unsolvable and highly uncertain queries without degrading their reasoning performance.
-We expect this work to advance the development of LLMs that understand not only how to solve problems, but also when not to attempt them.
-7 Future Work
-We plan to extend this work in three directions.
-First, we will construct diverse Out-of-Distribution (OOD) test sets to rigorously assess the generalization of refusal boundaries.
-Second, we aim to empirically investigate the root cause of Capability Collapse in smaller models, verifying our gradient interference hypothesis via feature space analysis.
-Finally, we will validate the scalability of our framework by applying it to a broader range of model architectures.
-Appendix A Appendix
-A.1 Theoretical Justification for False Unsolvability Penalty
-In this section, we provide a theoretical derivation to demonstrate why the penalty $\rho=-0.5$ is critical for preventing reward hacking, specifically the phenomenon of “collapse to universal rejection.”
-Problem Setup.
-Let $x$ be an input problem.
-Let $\mathcal{S}$ and $\mathcal{U}$ denote the sets of solvable and inherently unsolvable problems, respectively.
-Based on our dataset distribution, we assume balanced prior probabilities $P(x\in\mathcal{S})=P(x\in\mathcal{U})=0.5$.
-Let the model have a weak initial reasoning capability $\epsilon$ on solvable tasks (e.g., $\epsilon\approx 0.1$), meaning the probability of generating a correct solution given a solvable attempt is $P(\text{correct}\mid x\in\mathcal{S},\text{attempt})=\epsilon$.
-Conversely, identifying unsolvability is a binary decision: if the model predicts “unsolvable”, it receives a reward of +1+1 if $x\in\mathcal{U}$ and a penalty $\rho$ if $x\in\mathcal{S}$ (False Unsolvability).
-Decision Analysis.
-Consider the model’s decision process for a given input $x$ where the model has a posterior belief $p=P(x\in\mathcal{U})$ that the problem is unsolvable.
-The model compares the expected rewards $\mathbb{E}[R]$ of two actions: Attempting to Solve versus Rejecting.
-For the Attempt action, the model tries to generate a solution. The expected reward is the probability the problem is solvable multiplied by the success rate $\epsilon$:
-$\mathbb{E}[R_{\text{attempt}}]=(1-p)\cdot\epsilon\cdot(+1)+p\cdot 0=(1-p)\epsilon.$ (7)
-Note that attempting to solve an inherently unsolvable problem yields 0 reward (hallucination).
-For the Reject action, the model outputs “unsolvable”. The expected reward is the probability the problem is truly unsolvable plus the penalty for a false rejection:
-$\mathbb{E}[R_{\text{reject}}]=p\cdot(+1)+(1-p)\cdot\rho.$ (8)
-The Attempt Threshold.
-The model will choose to Attempt only when $\mathbb{E}[R_{\text{attempt}}]>\mathbb{E}[R_{\text{reject}}]$.
-Substituting the values, we get $(1-p)\epsilon>p+(1-p)\rho$.
-Solving for $p$, we derive the critical condition for attempting reasoning:
-$p<\frac{\epsilon-\rho}{1+\epsilon-\rho}.$ (9)
-This inequality defines the Attempt Threshold $T$. The model attempts reasoning only if its suspicion of unsolvability $p$ is lower than $T$.
-Case 1: No Penalty ($\rho=0$).
-Substituting $\epsilon=0.1$ and $\rho=0$, we get $T_{\text{zero}}=\frac{0.1}{1+0.1}\approx 0.09$.
-This implies the model will only attempt to solve if it is more than 91% sure the problem is solvable.
-Given the uncertainty in early training stages (where $p$ is often near 0.5), the condition $p<0.09$ is rarely met.
-The optimal strategy collapses to “Always Reject”, as it guarantees a risk-free average reward of 0.5 (since 50% of data is unsolvable) compared to the meager 0.05 expected from attempting.
-Case 2: With Penalty ($\rho=-0.5$).
-Substituting $\epsilon=0.1$ and $\rho=-0.5$, we get $T_{\text{penalty}}=\frac{0.1-(-0.5)}{1+0.1-(-0.5)}=\frac{0.6}{1.6}=0.375$.
-The threshold expands significantly from 0.09 to 0.375 (a $4\times$ increase).
-The model becomes willing to attempt reasoning even when it has moderate uncertainty about the problem’s solvability.
-This negative constraint effectively counteracts the low initial reasoning accuracy, forcing the model to explore solvable paths rather than collapsing into the local optimum of universal rejection.
-A.2 Theoretical Analysis of Capability-Refusal Dynamics
-In this section, we derive the mathematical relationship between the model’s reasoning capability and its tendency to refuse, explaining why a dynamic threshold is essential for preventing hallucination as the model becomes stronger.
-Decision Model.
-For a given solvable input $x$, let the model’s internal confidence (probability of generating the correct answer) be $p_{\text{correct}}$.
-The model chooses between two actions: Answer or Refuse.
-The expected reward for answering is:
-$\mathbb{E}[R_{\text{answer}}]=p_{\text{correct}}\cdot(+1)+(1-p_{\text{correct}})\cdot 0=p_{\text{correct}}$ (10)
-The expected reward for refusing is determined by the current batch accuracy $\beta$ and the target threshold $\tau$:
-$\mathbb{E}[R_{\text{refuse}}]=\lambda(\tau-\beta)$ (11)
-The model will choose to Refuse only if the expected reward of refusing exceeds that of answering:
-$\lambda(\tau-\beta)>p_{\text{correct}}\quad\implies\quad p_{\text{correct}}<\lambda(\tau-\beta)$ (12)
-Dynamics under Training.
-During the training process, the model’s reasoning capability improves, which leads to two simultaneous changes: 1.
-For a specific difficult problem, the model’s confidence $p_{\text{correct}}$ increases.
-2.
-The overall batch accuracy $\beta$ increases as the model solves more easy/medium problems correctly.
-Failure of Fixed Threshold (Fixed-$\tau$).
-If we fix the threshold $\tau$ to a constant value (e.g., $\tau=0.5$), as the model converges, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$).
-When $\beta>\tau$, the term $\lambda(\tau-\beta)$ becomes negative.
-Substituting this into Inequality 12, the condition becomes:
-$p_{\text{correct}}<\text{Negative Value}$ (13)
-Since probability $p_{\text{correct}}\in[0,1]$ is always non-negative, this condition is impossible to satisfy.
-Therefore, the model will never refuse, even if its confidence $p_{\text{correct}}$ is extremely low (e.g., 0.01).
-It is forced to hallucinate an answer to avoid the penalty of refusal.
-This explains the lower rejection rate and lower accuracy observed in the Fixed-$\tau$ experiments.
-Success of Dynamic Threshold.
-By dynamically setting $\tau$ (e.g., slightly higher than the moving average of $\beta$), we ensure that $(\tau-\beta)$ remains positive or near-zero.
-This maintains a valid decision boundary on the right side of Inequality 12, allowing the model to rationally refuse when its confidence $p_{\text{correct}}$ is truly low, regardless of how strong the model becomes overall.
-A.3 Theoretical Analysis of Capability Collapse in Small Models
-In this section, we propose a theoretical hypothesis for the capability collapse observed in smaller models (e.g., 1.7B) under the “No Data” setting.
-We frame this phenomenon as a potential instance of Gradient Interference arising from shared feature representations.
-Problem Definition.
-Let $x$ be an input problem.
-We define the model’s refusal policy via a scalar logit $h_{\theta}(x)=w^{\top}\phi(x)$, where $\phi(x)$ represents the features in the last hidden layer and $w$ is the weight vector of the refusal classification head.
-The probability of rejection is $P(\text{reject}|x)=\sigma(h_{\theta}(x))$.
-Training Dynamics under “No Data”.
-In the ablation setting, the training set $\mathcal{D}_{\text{train}}$ contains only solvable problems ($x_{s}\in\mathcal{S}$).
-The objective is to maximize the probability of answering (accepting) these problems.
-The loss function for a solvable sample $x_{s}$ can be formulated as minimizing the negative log-likelihood of not rejecting:
-$\mathcal{L}(\theta)=-\log(1-P(\text{reject}|x_{s}))=-\log(1-\sigma(w^{\top}\phi(x_{s})))$ (14)
-Computing the gradient with respect to $w$, we obtain:
-$\nabla_{w}\mathcal{L}=\frac{1}{1-\sigma}\cdot\frac{\partial(1-\sigma)}{\partial h}\cdot\phi(x_{s})=\sigma(w^{\top}\phi(x_{s}))\cdot\phi(x_{s})=P(\text{reject}|x_{s})\cdot\phi(x_{s})$ (15)
-Assuming simplified SGD with learning rate $\alpha$, the weight update is:
-$\Delta w\propto-\nabla_{w}\mathcal{L}=-\alpha\cdot P(\text{reject}|x_{s})\cdot\phi(x_{s})$ (16)
-Intuitively, whenever the model assigns a non-zero probability to rejecting a solvable problem (a false rejection), the optimizer pushes the weight vector $w$ in the direction opposite to $\phi(x_{s})$ to suppress this behavior.
-Hypothesized Impact on Unsolvable Test Data. Consider an unsolvable test instance $x_{u}\in\mathcal{U}$. The resulting change in its refusal logit $h(x_{u})$ would be:
-$\Delta h(x_{u})=(\Delta w)^{\top}\phi(x_{u})\propto-\alpha\cdot P(\text{reject}|x_{s})\cdot\left(\phi(x_{s})^{\top}\phi(x_{u})\right)$ (17)
-This derivation suggests that the impact on unsolvable instances depends critically on the geometric relationship between the feature representations of solvable and unsolvable problems:
-  • High Capacity Models (e.g., 4B+): We hypothesize that larger models can learn distinct, nearly orthogonal representations for solvable versus inherently contradictory problems (i.e., $\phi(x_{s})^{\top}\phi(x_{u})\approx 0$).
-In this case, suppressing rejection on $x_{s}$ has minimal impact on the model’s ability to reject $x_{u}$.
-  • Low Capacity Models (e.g., 1.7B): Smaller models may suffer from feature entanglement, relying on shared heuristics (such as sentence complexity or topic) to represent both problem types.
-If $\phi(x_{s})$ and $\phi(x_{u})$ are positively correlated ($\phi(x_{s})^{\top}\phi(x_{u})>0$), the term $\Delta h(x_{u})$ becomes negative.
-Conclusion.
-Under the assumption of feature entanglement in smaller models, training to answer solvable questions implicitly generates a negative gradient for the refusal logits of unsolvable questions.
-Without explicit negative samples ($x_{u}$) in the training set to provide a counter-gradient, this interference likely drives the refusal probability for unsolvable data toward zero, resulting in the observed capability collapse.
-A.4 Logic Puzzle Generation Algorithms
-A.4.1 Game24 Generation
+A.5 Logic Puzzle Generation Algorithms
+A.5.1 Game24 Generation
 The Game24 task involves combining a set of integers using arithmetic operations (+,,×,÷)(+,-$\times,\div$) to equal 24.
@@
 16:until Valid instance found
-A.4.2 Hamiltonian Cycle and Path Generation
+A.5.2 Hamiltonian Cycle and Path Generation
 The Hamiltonian Cycle problem asks whether a graph contains a closed loop visiting every vertex exactly once, while the Hamiltonian Path problem asks for a linear path visiting every vertex.
@@
 19:return $G$
-A.4.3 Hitori Generation
+A.5.3 Hitori Generation
 Hitori is a logic puzzle played on an $N\times N$ grid.
@@
 19:until Valid instance found
-A.4.4 Maze Navigation Generation
+A.5.4 Maze Navigation Generation
 The Maze Navigation task requires finding a valid path from a start position (S) to an end position (E) in a grid-based maze, where movement is restricted by walls.
@@
 To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement.
-We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction—a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths).
+We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction, a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths).
 We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS.
@@
 15:until Valid instance found
-A.5 Hyperparameters
-We implement our training pipeline using the verl (HybridFlow) framework (sheng2025hybridflow), which supports efficient reinforcement learning.
+A.6 Hyperparameters
+We implement our training pipeline using the verl (HybridFlow) framework and vllm engine (sheng2025hybridflow; kwon2023efficient), which supports efficient reinforcement learning.
 The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 6.
@@
 We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints.
+Specifically, the 1.7B model utilizes a larger batch size ($32\times 12$) compared to the 4B model ($14\times 12$).
+This adjustment balances the computational overhead, resulting in comparable training speeds for both architectures; typically, completing 120 training steps requires approximately one day on our 8 $\times$ NVIDIA A100 GPU cluster.
 Table 6: Hyperparameter settings used in our experiments.
@@
 Max Prompt Length 2048
+Generation Engine VLLM
 Infrastructure

### v3

@@
-Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems
+Learning the Boundary of Solvability: Aligning Large Language Models to Detect Unsolvable Problems
 Abstract
@@
 Second, we introduce UnsolvableRL, a reinforcement learning paradigm that balances objective unsolvability detection with calibrated confidence under capability limits.
-Empirically, our approach achieves near-perfect unsolvability detection ($90\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct.
+Empirically, our approach achieves robust unsolvability detection ($85\%$ detection rate) and boosts solvable reasoning accuracy from 43.4% to 69.4% on Qwen3-4B-Instruct.
 Crucially, we identify a data–training interaction: strict alignment constraints induce Capability Collapse without unsolvable data, but act as a regularizer for rigor when such data are included, thereby improving overall robustness.
 Our code and data are available at https://github.com/sfasfaffa/unsolvableQA.
-Learning the Boundary of Solvability: Aligning LLMs to Detect Unsolvable Problems
+Learning the Boundary of Solvability:
+Aligning Large Language Models to Detect Unsolvable Problems
 Dengyun Peng1,3thanks: Equal contribution   Qiguang Chen [^1] footnotemark: 1   Bofei Liu1   Jiannan Guan1   Libo Qin2thanks: Corresponding Author Zheng Yan1Jinhao Liu1Jianshu Zhang3Wanxiang Che122footnotemark: 2 1LARG, Research Center for Social Computing and Interactive Robotics, HIT 2School of Computer Science and Engineering, Central South University 3iFLYTEK
 1 Introduction
-As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (chen2025towards; wei2022chain; guo2025deepseek; jaech2024openai), a key research direction (qin2025largelanguagemodelsmeet) is ensuring they recognize their own limitations and behave reliably.
-Earlier work on confidence calibration (xiong2023can; li2024confidence; yoon2025reasoning; fisch2022calibrated) during reasoning underscores the need for models to know when not to answer.
-Further, studies on LLMs’ capability boundaries (chen2024unlocking) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; chen2025aware).
-More recently, TruthRL (wei2025truthrl) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
+As Large Language Models (LLMs) achieve strong performance on complex reasoning tasks (Chen et al., 2025c; Wei et al., 2022; Guo et al., 2025; Jaech et al., 2024; Hu et al., 2024), a key research direction (Qin et al., 2025) is ensuring they recognize their own limitations and behave reliably.
+Earlier work on confidence calibration (Xiong et al., 2023; Li et al., 2024; Yoon et al., 2025; Fisch et al., 2022) during reasoning underscores the need for models to know when not to answer.
+Further, studies on LLMs’ capability boundaries (Chen et al., 2024) examine the self-awareness of their limitations to avoid unproductive reasoning on problems they are likely to fail (Zhang et al., 2025; Mei et al., 2025; Kalai et al., 2025; Chen et al., 2025b).
+More recently, TruthRL (Wei et al., 2025) has encouraged LLMs to refuse when uncertain through a ternary reward mechanism.
 Figure 1: Comparison of alignment paradigms.
@@
 First, to support rigorous analysis of unsolvability, we develop a fully automated synthesis pipeline and introduce UnsolvableQA, a dataset of paired solvable and unsolvable instances spanning puzzles and mathematics.
-For programmatically verifiable puzzle tasks, we construct provably unsolvable instances by injecting code-level logical contradictions.
-For mathematics problems without reliable programmatic verification, inspired by lightman2024lets; ling2023deductive, we propose a “Reverse Construction” method: starting from a solvable problem and its correct rationale, we introduce a controlled contradiction into the rationale and then prompt LLMs to generate an unsolvable problem that matches the resulting flawed reasoning.
-Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (shao2024deepseekmath).
+For mathematics, inspired by Lightman et al.
+(2024); Ling et al.
+(2023), we propose a “Reverse Construction” method that injects controlled contradictions into valid rationales to generate matching unsolvable problems.
+For programmatically verifiable puzzles, we construct instances via code-level logical contradictions.
+Notably, our logic puzzle pipeline enables the rapid generation of infinite instances with arbitrary difficulty, and we will fully release our construction code upon acceptance.
+Then, to train models that remain robust on both solvable and unsolvable inputs, we propose UnsolvableRL, a reinforcement-learning (RL) method based on GRPO (Shao et al., 2024).
 Unlike prior approaches that conflate refusal types, we define 3 distinct RL objectives: (1) Inherent unsolvability rewards generating an explicit refusal tag when solvability is detected; (2) False unsolvability Penalty discourages erroneous refusals by penalizing labeling solvable problems as unsolvable, thereby requiring explicit solvability checks before refusal; (3) Capability calibration uses a mechanism that rewards refusal only when the problem difficulty exceeds the LLMs’ current capability.
-Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 40.3% to 90.7% (near-perfect) and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 80.1%.
+Empirical results show a synergistic effect: UnsolvableRL elevates the unsolvable problem detection rate of Qwen3-4B from 38.8% to 87.5% and boosts its solvable problem accuracy from 43.4% to 69.4%, achieving an overall score of 78.6%.
 Our ablation study reveals three key insights: (1) models cannot learn to detect unsolvable problems without explicit unsolvable training data; (2) penalizing false refusals without such data causes “capability collapse”, where models stop refusing altogether; (3) when paired with unsolvable data, the penalty acts as a regularizer, trading a small drop in detection for stronger reasoning accuracy, a Pareto improvement over baselines.
@@
 Automatic Reasoning Dataset Construction.
-As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (liu2024automatic; wang2023data).
+As the reliance of Large Language Models on data quality increases, Automated Dataset Construction (ADC) has emerged as a prominent research direction (Liu et al., 2024; Wang et al., 2023).
 In the realm of reasoning, recent efforts on automatic data construction primarily focus on ensuring data solvability and verifiability through formal or programmatic means.
-Representative works such as Enigmata (chen2025enigmata) and BIGGSM (chen2024unlocking) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively.
-Concurrently, research on verifying the correctness of rationale traces has also been advancing (ling2023deductive; chen2025towards).
+Representative works such as Enigmata (Chen et al., 2025a) and BIGGSM (Chen et al., 2024) have significantly enriched the resources for solvable reasoning tasks by employing synthetic puzzles and difficulty-graded annotations, respectively.
+Concurrently, research on verifying the correctness of rationale traces has also been advancing (Ling et al., 2023; Chen et al., 2025c).
 In contrast, progress in dataset construction regarding problem unsolvability has been relatively slow.
-Earlier, SQuAD 2.0 (rajpurkar2018knowdontknowunanswerable) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings.
+Earlier, SQuAD 2.0 (Rajpurkar et al., 2018) examined whether questions are answerable given a long context, but it does not adequately evaluate refusal behavior in complex logical-reasoning settings.
 Consequently, for reasoning tasks, existing data-synthesis efforts focus mainly on solvable instances while largely neglecting unsolvable questions, leaving a significant research gap.
 Figure 2: The UnsolvableQA data collection pipeline.
-(a) Mathematical Data Collection: We employ a “Reverse Construction” method where contradictions are injected into the solution path of solvable problems.
+(a) Mathematical Data Collection: It employs a “Reverse Construction” method where contradictions are injected into the solution path of solvable problems.
 The resulting instances are verified by an LLM to ensure they lead to a contradiction (unsolvable) rather than a valid solution.
-(b) Puzzle Data Collection: We utilize domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
+(b) Puzzle Data Collection: It utilizes domain-specific programmatic generators and automated deterministic verifiers (e.g., SAT solvers, DFS) to rigorously classify randomly generated instances into solvable and unsolvable sets.
 Reinforcement Learning for Reasoning.
-Beyond dataset construction, Reinforcement Learning (RL) (ouyang2022traininglanguagemodelsfollow; schulman2017proximalpolicyoptimizationalgorithms; rafailov2024directpreferenceoptimizationlanguage) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness.
-Recent RL algorithms for reasoning, such as GRPO (shao2024deepseekmath) and DAPO (yu2025dapoopensourcellmreinforcement), further enhance models’ reasoning abilities (sheng2025hybridflow).
-Concurrent work also leverages RL to improve reasoning efficiency under cost constraints (tu2025learning; team2025kimi; guo2025deepseek; wang2025aspoasymmetricimportancesampling; chen2025aware).
-In parallel, advances such as TruthRL (wei2025truthrl) introduce objectives that encourage models to refuse likely hallucinations, and research on self-awareness enables models to decline answering when appropriate (zhang2025self; mei2025reasoning; kalai2025languagemodelshallucinate; Steyvers2025).
+Beyond dataset construction, Reinforcement Learning (RL) (Ouyang et al., 2022; Schulman et al., 2017; Rafailov et al., 2024) has become a central approach for aligning LLMs with desired behaviors such as safety and truthfulness.
+Recent RL algorithms for reasoning, such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), further enhance models’ reasoning abilities (Sheng et al., 2025).
+Concurrent work also leverages RL to improve reasoning efficiency under cost constraints (Tu et al., 2025; Team et al., 2025; Guo et al., 2025; Wang et al., 2025; Chen et al., 2025b).
+In parallel, advances such as TruthRL (Wei et al., 2025) introduce objectives that encourage models to refuse likely hallucinations, and research on self-awareness enables models to decline answering when appropriate (Zhang et al., 2025; Mei et al., 2025; Kalai et al., 2025; Steyvers et al., 2025).
 Despite these advancements, existing frameworks largely conflate objective unsolvability with subjective uncertainty.
@@
 The model is then prompted to reassess the problem’s solvability given this complete contextual guidance.
-Only instances that are ultimately verified as unsolvable through this two-tier procedure are retained.
+Only instances verified as unsolvable are retained, and we manually verified the final test set.
 3.2 Logic Puzzle Generation
@@
 While training prompts follow templated formats and pool Maze difficulties, test prompts are diversified and preserve fine-grained splits (including distinct AIME years [^1] AIME year ranges: train 1983–2023; test 2024–2025.).
-All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification.
+All instances undergo rigorous quality control: logic puzzles are verified via SAT/CSP/graph solvers, mazes via deterministic search, and math problems through multi-pass LLM verification and manual checks.
 We also introduce two OOD testsets in Appendix A.2.
@@
 Specifically, to satisfy three distinct behavioral objectives, our training framework optimizes a policy $\pi_{\theta}$ as follows.
-First, for solvable and easy instances within the model’s capabilities ($x\in\mathcal{S}\cap\mathcal{E}$), the objective is Reasoning Accuracy, requiring the model to produce correct reasoning chains and final answers.
+First, for solvable and subjectively easy instances within the model’s capabilities ($x\in\mathcal{S}\cap\mathcal{E}$), the objective is Reasoning Accuracy, requiring the model to produce correct reasoning chains and final answers.
 Second, for inherently unsolvable instances ($x\in\mathcal{U}\cap\mathcal{E}$), the objective is Unsolvability Detection: the model must explicitly recognize the inconsistency and output a dedicated tag (e.g., <unsolvable>) instead of hallucinating a solution, which we treat as an objective correctness task.
-Third, for hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
+Third, for subjectively hard yet solvable instances ($x\in\mathcal{H}$) that exceed the model’s effective reasoning capacity (as indicated by low internal confidence), the objective is Capability Calibration, where the model should prudently abstain (e.g., output <beyond_capacity>) to reduce false positives.
 4.2 Group-Relative Policy Optimization
@@
 where $\mu_{\mathbf{R}}$ and $\sigma_{\mathbf{R}}$ are the mean and standard deviation of the rewards $\mathbf{R}=\{R(y_{j}\mid x)\}_{j=1}^{G}$ within the group.
-This group-based normalization can stabilize training due to reward-scale differences across solvable and unsolvable domains.
+This group-based reward facilitates efficient and stable training, helps effectively distinguish different responses to the same problem, and eliminates the need for critic and reward models that occupy massive GPU memory.
 4.3 Reward Design
 SOTA Baselines Qwen3-1.7B Qwen3-4B
-Dataset Metric Deepseek GPT-5.1 Gemini-3 Instruct + UnsolRL Instruct + UnsolRL
+Dataset Metric Deepseek-V3.2-R GPT-5.1-Low Gemini-3 Instruct + UnsolvableRL Instruct + UnsolvableRL
 Game24 S / U 98.0 / 77.5 14.0 / 50.0 94.0 / 94.0 78.0 / 23.0 84.0 / 100.0 81.0 / 17.0 94.5 / 99.0
-Mean 87.8 32.0 94.0 50.5 92.0 49.0 96.8
+Mean 87.8 32.0 94.0 50.5 92.0 (+41.5) 49.0 96.8 (+47.8)
 HamCycle S / U 83.5 / 94.0 8.3 / 82.0 100.0 / 98.0 22.9 / 28.0 20.8 / 88.0 37.5 / 57.0 41.1 / 94.5
-Mean 88.8 45.1 99.0 25.5 54.4 47.3 67.8
+Mean 88.8 45.1 99.0 25.5 54.4 (+28.9) 47.3 67.8 (+20.5)
 HamPath S / U 84.0 / 96.0 50.0 / 86.0 100.0 / 90.0 14.0 / 33.0 27.0 / 85.0 37.0 / 56.5 55.0 / 96.5
-Mean 90.0 68.0 95.0 23.5 56.0 46.8 75.8
+Mean 90.0 68.0 95.0 23.5 56.0 (+32.5) 46.8 75.8 (+29.0)
 Hitori S / U 70.0 / 86.0 6.0 / 12.0 98.0 / 100.0 8.0 / 21.0 11.0 / 59.0 34.5 / 6.5 63.5 / 94.5
-Mean 78.0 9.0 99.0 14.5 35.0 20.5 79.0
+Mean 78.0 9.0 99.0 14.5 35.0 (+20.5) 20.5 79.0 (+58.5)
 Maze (Easy) S / U 100.0 / 98.9 95.0 / 100.0 99.0 / 94.6 0.0 / 85.6 0.0 / 95.2 32.7 / 55.6 96.5 / 98.9
-Mean 99.5 97.5 96.8 42.8 47.6 44.1 97.7
+Mean 99.5 97.5 96.8 42.8 47.6 (+4.8) 44.1 97.7 (+53.6)
 Maze (Hard) S / U 98.0 / 98.0 86.0 / 94.0 96.0 / 91.0 0.0 / 80.0 0.0 / 79.0 13.3 / 64.5 65.8 / 89.0
-Mean 98.0 90.0 93.5 40.0 39.5 38.9 77.9
-AIME 24-25 S / U 85.0 / 63.8 61.7 / 48.9 95.0 / 34.1 38.3 / 33.5 35.4 / 52.7 67.9 / 25.3 69.6 / 62.2
-Mean 74.4 55.3 64.5 35.9 44.1 46.6 65.9
-Overall S / U 88.4 / 87.7 45.9 / 67.6 97.4 / 86.0 23.0 / 43.4 25.5 / 79.8 43.4 / 40.3 69.4 / 90.7
-Mean 88.1 56.7 91.7 33.2 52.7 41.9 80.1
+Mean 98.0 90.0 93.5 40.0 39.5 (-0.5) 38.9 77.9 (+39.0)
+AIME 24-25 S / U 85.0 /40.3 61.7 / 42.5 95.0 / 21.2 38.3 / 21.2 35.4 / 28.7 67.9 / 14.8 69.6 / 40.4
+Mean 62.7 52.1 58.1 29.8 32.1 (+2.3) 41.4 55.0 (+13.6)
+Overall S / U 88.4 / 84.4 45.9 / 66.6 97.4 / 84.1 23.0 / 41.7 25.5 / 76.4 43.4 / 38.8 69.4 / 87.5
+Mean 86.1 56.2 90.8 32.4 50.9(+18.5) 41.1 78.6 (+37.5)
 Table 2: Detailed performance comparison.
 For each dataset, the first row reports Solvable Accuracy (S) / Unsolvable Detection Rate (U), and the second row reports their Mean (M).
-We bold the global best score across all models for each metric.
-Qwen3-4B + UnsolRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
+Values in parentheses indicate the absolute improvement over the base Instruct model.
+We bold the global best score across all models.
+Qwen3-4B + UnsolvableRL achieves SOTA performance in Unsolvable Detection (Overall U) and Game24 Mean accuracy.
 As illustrated in Figure 3, our framework explicitly distinguishes between three types of model responses: (1) Reward for Reasoning Accuracy (providing a solution), (2) Reward for Unsolvability Detection (identifying inherent contradictions), and (3) Reward for Capability Calibration (admitting capability limits).
@@
 (1) A Correct Response (+1+1) is assigned when the model outputs the correct answer for a solvable instance.
-(2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward.
-This asymmetric scheme emphasizes reasoning capability by rewarding successful problem solving while not explicitly penalizing unsuccessful but genuine reasoning attempts on solvable tasks.
+(2) For Incorrect Reasoning (0) on solvable instances, such as producing an incorrect final answer despite performing the calculation, the model receives a neutral reward (0).
+This ensures that honest calculation errors are not penalized, maintaining the model’s incentive to attempt reasoning.
 Reward for Unsolvability Detection ($R_{\text{detect}}$).
@@
 (2) To prevent degenerate behavior, we impose a False Unsolvability Penalty ($\rho=-0.5$): if the model incorrectly declares a solvable problem unsolvable, it receives a negative reward.
-As derived in our decision-theoretic analysis in Appendix A.3, this penalty is mathematically critical to prevent the model from collapsing into a local optimum of universal rejection.
-It effectively expands the decision threshold, encouraging the model to attempt reasoning even under moderate uncertainty.
+This penalty effectively expands the decision threshold, encouraging the model to attempt reasoning even under moderate uncertainty.
+As derived in our decision-theoretic analysis in Appendix A.5, this term is mathematically critical to prevent the model from collapsing into a local optimum of universal rejection.
+Moreover, our empirical ablation on $\rho$ selection identifies $\rho=-0.5$ as the optimal value; while stricter penalties yield marginal gains in solvable accuracy, they significantly impair unsolvable detection capabilities, making this value the most effective trade-off.
+We refer to this penalty mechanism as $P$ in our subsequent experimental analysis and tables.
 (3) For Error-Detect Reasoning (0) on unsolvable instances, such as hallucinating a solution instead of outputting unsolvable, the model receives a neutral reward.
@@
 $R=R_{\text{acc}}+R_{\text{detect}}+R_{\text{cal}}.$ (5)
+Figure 4: Detailed Ablation Metrics.
+Breakdown of Solvable Accuracy (S, Red), Unsolvable Detection Rate (U, Blue), and Mean Performance (M, Green) across different configurations.
+The w/o UnsData (Unsolvable Data) w/ P setting shows a catastrophic drop in U for the 1.7B model (4.2%), indicating capability collapse.
+Notably, w/ UnsData w/o P achieves the highest Mean (M) score by maximizing U.
+The full method (w/ UnsData w/ P) demonstrates a valuable trade-off, boosting S while maintaining a high U, optimizing for practical scenarios.
 4.4 Answer Extraction and Verification
@@
 5 Experiments
-Figure 4: Detailed Ablation Metrics.
-Breakdown of Solvable Accuracy (S, Red), Unsolvable Detection Rate (U, Blue), and Mean Performance (M, Green) across different configurations.
-The w/o UnsData w/ P setting shows a catastrophic drop in U for the 1.7B model (4.2%), indicating capability collapse.
-Notably, w/ UnsData w/o P achieves the highest Mean (M) score by maximizing U.
-The full method (w/ UnsData w/ P) demonstrates a valuable trade-off, boosting S while maintaining a high U, optimizing for practical scenarios.
-Figure 5: Mechanism Analysis: Trade-off between $S$ and $U$.
-Visualizing the interaction between Data and Penalty.
-The Purple Dashed Arrow highlights the comprehensive enhancement: compared to the naive baseline (Red Square), our full method (Blue Circle) achieves significant improvements in both Solvable Accuracy ($S$) and Unsolvable Detection ($U$).
-Specific dynamics include: (left) Capability Collapse: Applying the penalty without data causes the 1.7B model to collapse (Red Arrow).
-(right) Strategic Trade-off: With adequate data, the penalty acts as a regularizer (Blue Arrow), trading a small amount of $U$ for a gain in $S$ to optimize for real-world applications.
-We trained and evaluated Qwen3-1.7B and Qwen3-4B (yang2025qwen3technicalreport) in think mode on UnsolvableQA.
+Figure 5: Visualizing the interaction between Solvable Accuracy ($S$) and Unsolvable Detection ($U$).
+The Purple Arrow highlights our full method (Blue Circle) achieving a Pareto improvement over the baseline (Red Square).
+Both models exhibit consistent dynamics: applying penalties without Unsolvable Data (UD) triggers Capability Collapse (Red Arrow), sacrificing $U$ to maximize $S$.
+Conversely, when paired with UD, the penalty acts as a Regularizer (Blue Arrow), trading marginal $U$ for significant gains in $S$ compared to using UD alone (Blue Square).
+We trained and evaluated Qwen3-1.7B and Qwen3-4B (Yang et al., 2025) in think mode on UnsolvableQA.
 For evaluation, we report average@8 on the AIME partitions and average@4 on other datasets.
 We standardize training steps: ablation variants are trained for 120 steps, while the final models are trained for 320 steps.
-5.1 Main Results: Bridging the Gap with SOTA
-Table 2 compares our Qwen3-UnsolRL models against both proprietary SOTA models and base baselines.
-The complete experimental results are presented in Appendix 4.
-High-reasoning LLMs (e.g., Deepseek-V3.2-R1) inherently perform well, whereas base models struggle significantly (e.g., Qwen3-1.7B Instruct achieves only 33.2% mean score).
-Our method bridges this gap: Qwen3-4B + UnsolRL achieves a global mean score of 80.1%, significantly outperforming GPT-5.1-low (56.7%) and approaching the performance of Gemini-3 (91.7%).
-Notably, our 4B model achieves the highest global Unsolvable Detection rate (90.7%), surpassing even the much larger Deepseek and Gemini models.
+5.1 Main Results: Comparative Analysis with SOTA Models
+Table 2 compares our Qwen3-UnsolvableRL models against both proprietary SOTA models and base baselines.
+The complete experimental results are presented in Appendix 4, which also verifies the generalizability of our method on Llama models, where UnsolvableRL boosts the overall Mean score to 55.1% compared to the 25.2% baseline.
+High-reasoning LLMs (e.g., Deepseek-V3.2-R1) inherently perform well, whereas base models struggle significantly (e.g., Qwen3-1.7B Instruct achieves only 32.4% mean score).
+Our method bridges this gap: Qwen3-4B + UnsolvableRL achieves a global mean score of 78.6%, significantly outperforming GPT-5.1-low (56.2%) and approaching Gemini-3 (90.8%).
+Notably, our 4B model achieves the highest global Unsolvable Detection rate (87.5%), surpassing even the much larger Deepseek and Gemini models.
 This confirms that equipping models with the ability to identify unsolvability ($\mathcal{U}$) strengthens the rigor required to solve feasible problems ($\mathcal{S}$).
+We further validated our method on two OOD benchmarks in Appendix A.2.
 Figure 6: Performance Comparison and Calibration Analysis.
@@
 As shown in Figure 4, the naive baseline trained without unsolvable data (w/o UnsData) lacks the fundamental capability to recognize unsolvability.
-For Qwen3-4B, detection is weak ($U=64.1\%$).
-Introducing explicit unsolvable data provides the necessary supervision, boosting detection significantly ($U\to 89.3\%$).
-This confirms that instruction following alone is insufficient; explicit data is the foundational requirement for boundary learning.
+For Qwen3-4B, detection is weak ($U=62.6\%$).
+Introducing explicit unsolvable data provides the necessary supervision, boosting detection significantly ($U\to 84.8\%$).
+This confirms that optimizing solely for reasoning correctness on solvable tasks is insufficient.
 Applying Penalty without Data Triggers Capability Collapse.
@@
 When the penalty is enforced without providing ground-truth unsolvable examples, the model, lacking positive examples of valid refusal, minimizes the penalty risk by defaulting to answering all queries.
-This behavior effectively collapses its refusal mechanism, as seen in the 1.7B model’s detection rate plummeting from 70.8% to 4.2%.
+This behavior effectively collapses its refusal mechanism, as seen in the 1.7B model’s detection rate plummeting from 68.8% to 4.2%.
 Applying Penalty with Data Facilitates a Strategic Trade-off.
 When grounded with adequate data, the penalty shifts from being a suppressor to a regularizer.
-For Qwen3-4B, adding the penalty (w/ UnsData w/ P) improves Solvable Accuracy ($S$) from 60.6% to 62.6%, at the cost of a slight drop in detection ($U:89.3\%\to 84.8\%$).
+For Qwen3-4B, adding the penalty (w/ UnsData w/ P) improves Solvable Accuracy ($S$) from 60.6% to 62.6%, at the cost of a slight drop in detection ($U:84.8\%\to 81.5\%$).
 This trade-off is strategic: it sacrifices marginal detection performance to enforce stricter reasoning rigor on feasible problems, optimizing the model for real-world utility.
@@
 Viewing the trajectory globally (the purple dashed arrow in Figure 5b), our full method represents a comprehensive upgrade over the naive baseline (w/o UnsData w/o P).
-We achieve significant gains in both dimensions for Qwen3-4B ($S:57.7\%\to 62.6\%$, $U:64.1\%\to 84.8\%$).
+We achieve significant gains in both dimensions for Qwen3-4B ($S:57.7\%\to 62.6\%$, $U:62.6\%\to 81.5\%$).
 This confirms that UnsolvableRL does not merely shift the decision boundary, but fundamentally enhances both reliability and reasoning capability.
 5.3 Comparative Analysis: TruthRL and Calibration
-We further benchmark our method against TruthRL(wei2025truthrl) and different thresholding strategies. The results are summarized in Figure 6.
+We further benchmark our method against TruthRL (Wei et al., 2025) and different thresholding strategies.
 Superior Balance Compared to TruthRL.
 As shown in the first two panels of Figure 6, TruthRL tends to be overly conservative, leading to over-refusal.
-For Qwen3-1.7B, while TruthRL achieves a high Rejection Rate (18.58%), its Accuracy drops to 40.4%, which is inferior to even the baseline.
-In contrast, our method achieves a significantly higher Accuracy (50.3%) with a moderate and precise Rejection Rate (7.96%).
-This indicates that our False Unsolvability Penalty provides a softer, more distinctive signal than TruthRL’s hard refusal reward, preserving the model’s core reasoning capabilities.
+For Qwen3-1.7B, while TruthRL achieves a high Rejection Rate (18.58%), its Accuracy drops to 39.0%, which is inferior to even the baseline.
+In contrast, our method achieves a significantly higher Accuracy (45.5%) with a moderate and precise Rejection Rate (10.58%).
+This indicates that our target accuracy threshold provides a softer, more distinctive signal than TruthRL’s hard refusal reward, preserving the model’s core reasoning capabilities.
 Dynamic Thresholding Prevents Overconfidence.
@@
 The Fixed-$\tau$ baseline yields a low rejection rate (4.36%), suggesting overconfidence.
-By dynamically adjusting $\tau$ based on the moving average of solvable confidence, our method (w/ P Final) maintains a robust rejection rate (6.87%) without sacrificing accuracy (80.1%), ensuring the model remains prudent even as it becomes stronger.
+By dynamically adjusting $\tau$ based on the moving average of solvable confidence, our method (w/ P Final) maintains a robust rejection rate (6.87%) without sacrificing accuracy (78.6%), ensuring the model remains prudent even as it becomes stronger.
 6 Conclusion
@@
 We introduce UnsolvableQA for rigorous evaluation and UnsolvableRL, an alignment framework featuring dynamic calibration.
-Empirically, our method achieves a Pareto improvement over baselines: it boosts solvable accuracy (Overall $S$) from 43.4% to 69.4% while maintaining robust unsolvability detection ($U>90\%$ on Qwen3-4B).
+Empirically, our method achieves a Pareto improvement over baselines: it boosts solvable accuracy (Overall $S$) from 43.4% to 69.4% while maintaining robust unsolvability detection ($U>85\%$ on Qwen3-4B).
 Crucially, our analysis uncovers the dual nature of alignment penalties: without grounding data, they trigger Capability Collapse; yet with adequate data, they facilitate a Strategic Trade-off, acting as a Regularizer for Rigor that prevents hallucination without the over-refusal pitfalls observed in prior methods like TruthRL.
@@
 Future work could expand the dataset to include more diverse and ambiguous types of unsolvable problems.
+Additionally, due to computational resource constraints, we were unable to conduct a more exhaustive set of ablation studies, such as scaling experiments on larger model architectures.
 Appendix A Appendix
@@
 Checkpoint S U M S U M S U M S U M S U M S U M S U M $\mathrm{S}$ $\mathrm{U}$ $\mathrm{M}$
-Deepseek-V3.2-R 98.0 77.5 87.8 83.5 94.0 88.8 84.0 96.0 90.0 70.0 86.0 78.0 100.0 98.9 99.5 98.0 98.0 98.0 85.0 63.8 74.4 88.4 87.7 88.1
-GPT-5.1-Low 14.0 50.0 32.0 8.3 82.0 45.1 50.0 86.0 68.0 6.0 12.0 9.0 95.0 100.0 97.5 86.0 94.0 90.0 61.7 48.9 55.3 45.9 67.6 56.7
-Gemini-3.0-Pro 94.0 94.0 94.0 100.0 98.0 99.0 100.0 90.0 95.0 98.0 100.0 99.0 99.0 94.6 96.8 96.0 91.0 93.5 95.0 34.1 64.5 97.4 86.0 91.7
-Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 13.3 64.5 38.9 67.9 25.3 46.6 43.4 40.3 41.9
-+ TruthRL 91.0 91.0 91.0 32.3 86.0 59.2 49.0 90.0 69.5 38.0 15.0 26.5 90.0 77.1 83.6 31.0 56.5 43.8 67.5 25.5 46.5 57.0 63.0 60.0
-+ w/o UnsData w/o P 83.0 92.0 87.5 36.5 81.0 58.8 47.0 84.0 65.5 46.0 20.0 33.0 88.5 76.6 82.6 34.5 60.5 47.5 68.3 34.6 51.5 57.7 64.1 60.9
-+ w/o UnsData w/ P 90.5 12.5 51.5 39.6 67.5 53.5 49.5 64.5 57.0 56.5 4.0 30.3 94.8 83.2 89.0 56.3 72.0 64.1 69.0 18.4 43.7 65.2 46.0 55.6
-+ w/ UnsData w/o P 84.0 95.0 89.5 39.6 96.0 67.8 53.0 100.0 76.5 45.0 87.0 66.0 95.5 95.2 95.4 40.0 89.5 64.8 67.1 62.2 64.7 60.6 89.3 75.0
-+ w/ UnsData w/ P 88.0 97.0 92.5 35.9 95.5 65.7 51.5 99.5 75.5 48.5 71.0 59.8 94.0 95.5 94.8 53.3 79.8 66.6 66.9 55.3 61.1 62.6 84.8 73.7
-+ Fixed-$\tau$ Final 87.5 99.0 93.3 39.1 100.0 69.6 67.0 99.5 83.3 51.0 98.0 74.5 96.3 100.0 98.2 52.5 95.8 74.2 69.0 63.6 66.3 66.1 93.7 79.9
-+ w/ UnsData w/ P Final 94.5 99.0 96.8 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.5 98.9 97.7 65.8 89.0 77.9 69.6 62.2 65.9 69.4 90.7 80.1
-Qwen3-1.7B-Instruct 78.0 23.0 50.5 22.9 28.0 25.5 14.0 33.0 23.5 8.0 21.0 14.5 0.0 85.6 42.8 0.0 80.0 40.0 38.3 33.5 35.9 23.0 43.4 33.2
-+ TruthRL 76.0 80.0 78.0 19.8 49.0 34.4 17.0 58.0 37.5 4.0 25.0 14.5 1.5 82.4 42.0 0.0 73.5 36.8 42.1 37.2 39.7 22.9 57.9 40.4
-+ w/o UnsData w/o P 77.0 88.0 82.5 25.0 60.0 42.5 15.0 83.0 49.0 5.0 53.0 29.0 0.5 93.6 47.1 0.0 79.5 39.8 32.9 38.3 35.6 22.2 70.8 46.5
+Deepseek-V3.2-R 98.0 77.5 87.8 83.5 94.0 88.8 84.0 96.0 90.0 70.0 86.0 78.0 100.0 98.9 99.5 98.0 98.0 98.0 85.0 40.3 62.7 88.4 84.4 86.5
+GPT-5.1-Low 14.0 50.0 32.0 8.3 82.0 45.1 50.0 86.0 68.0 6.0 12.0 9.0 95.0 100.0 97.5 86.0 94.0 90.0 61.7 42.5 52.1 45.9 66.6 56.2
+Gemini-3.0-Pro 94.0 94.0 94.0 100.0 98.0 99.0 100.0 90.0 95.0 98.0 100.0 99.0 99.0 94.6 96.8 96.0 91.0 93.5 95.0 21.2 58.1 97.4 84.1 90.8
+Qwen3-4B-Instruct 81.0 17.0 49.0 37.5 57.0 47.3 37.0 56.5 46.8 34.5 6.5 20.5 32.7 55.6 44.1 13.3 64.5 38.9 67.9 14.8 41.4 43.4 38.8 41.1
++ TruthRL 91.0 91.0 91.0 32.3 86.0 59.2 49.0 90.0 69.5 38.0 15.0 26.5 90.0 77.1 83.6 31.0 56.5 43.8 67.5 19.1 43.3 57.0 62.1 59.6
++ w/o UnsData w/o P 83.0 92.0 87.5 36.5 81.0 58.8 47.0 84.0 65.5 46.0 20.0 33.0 88.5 76.6 82.6 34.5 60.5 47.5 68.3 24.0 46.2 57.7 62.6 60.2
++ w/o UnsData w/ P 90.5 12.5 51.5 39.6 67.5 53.5 49.5 64.5 57.0 56.5 4.0 30.3 94.8 83.2 89.0 56.3 72.0 64.1 69.0 7.4 38.2 65.2 44.4 54.8
++ w/ UnsData w/o P 84.0 95.0 89.5 39.6 96.0 67.8 53.0 100.0 76.5 45.0 87.0 66.0 95.5 95.2 95.4 40.0 89.5 64.8 67.1 30.8 49.0 60.6 84.8 72.7
++ w/ UnsData w/ P 88.0 97.0 92.5 35.9 95.5 65.7 51.5 99.5 75.5 48.5 71.0 59.8 94.0 95.5 94.8 53.3 79.8 66.6 66.9 31.9 49.4 62.6 81.5 72.0
++ Fixed-$\tau$ Final 87.5 99.0 93.3 39.1 100.0 69.6 67.0 99.5 83.3 51.0 98.0 74.5 96.3 100.0 98.2 52.5 95.8 74.2 69.0 41.9 55.5 66.1 90.6 78.4
++ w/ UnsData w/ P Final 94.5 99.0 96.8 41.1 94.5 67.8 55.0 96.5 75.8 63.5 94.5 79.0 96.5 98.9 97.7 65.8 89.0 77.9 69.6 40.4 55.0 69.4 87.5 78.6
+Qwen3-1.7B-Instruct 78.0 23.0 50.5 22.9 28.0 25.5 14.0 33.0 23.5 8.0 21.0 14.5 0.0 85.6 42.8 0.0 80.0 40.0 38.3 21.2 29.8 23.0 41.7 32.4
++ TruthRL 76.0 80.0 78.0 19.8 49.0 34.4 17.0 58.0 37.5 4.0 25.0 14.5 1.5 82.4 42.0 0.0 73.5 36.8 42.1 17.0 29.6 22.9 55.0 39.0
++ w/o UnsData w/o P 77.0 88.0 82.5 25.0 60.0 42.5 15.0 83.0 49.0 5.0 53.0 29.0 0.5 93.6 47.1 0.0 79.5 39.8 32.9 24.4 28.7 22.2 68.8 45.5
 + w/o UnsData w/ P 85.0 2.0 43.5 22.4 10.5 16.5 25.5 14.0 19.8 12.0 2.5 7.3 0.0 0.0 0.0 0.0 0.0 0.0 40.8 0.5 20.7 26.5 4.2 15.4
-+ w/ UnsData w/o P 70.0 97.0 83.5 22.9 94.0 58.5 22.0 97.0 59.5 4.0 79.0 41.5 0.5 95.2 47.9 0.0 86.0 43.0 31.7 56.4 44.0 21.6 86.4 54.0
-+ w/ UnsData w/ P 82.0 96.5 89.3 18.2 85.0 51.6 20.5 89.5 55.0 8.0 62.0 35.0 0.0 90.7 45.4 0.0 74.3 37.2 40.8 36.4 38.6 24.2 76.3 50.3
-+ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2
++ w/ UnsData w/o P 70.0 97.0 83.5 22.9 94.0 58.5 22.0 97.0 59.5 4.0 79.0 41.5 0.5 95.2 47.9 0.0 86.0 43.0 31.7 46.8 39.3 21.6 85.0 53.3
++ w/ UnsData w/ P 82.0 96.5 89.3 18.2 85.0 51.6 20.5 89.5 55.0 8.0 62.0 35.0 0.0 90.7 45.4 0.0 74.3 37.2 40.8 21.2 31.0 24.2 74.2 49.2
++ w/ UnsData w/ P Final 84.0 100.0 92.0 20.8 88.0 54.4 27.0 85.0 56.0 11.0 59.0 35.0 0.0 95.2 47.6 0.0 79.0 39.5 35.4 28.7 32.1 25.5 76.4 50.9
+Llama3-8B-Distill 28.0 48.0 38.0 27.1 2.0 14.6 14.0 0.0 7.0 6.0 56.0 31.0 1.0 44.7 22.9 0.0 52.0 26.0 35.0 39.4 37.2 15.9 34.6 25.2
++ w/o UnsData w/o P 58.0 2.0 30.0 41.7 0.0 20.8 50.0 0.0 25.0 6.0 2.0 4.0 67.0 7.4 37.2 8.0 8.0 8.0 46.7 9.6 28.1 39.6 4.1 21.9
++ w/ UnsData w/ P 66.0 4.0 35.0 39.6 98.0 68.8 48.0 90.0 69.0 10.0 78.0 44.0 60.0 3.2 31.6 7.0 20.0 13.5 50.0 30.9 40.5 40.1 46.3 43.2
++ w/ UnsData w/ P Final 80.0 22.0 51.0 52.1 100.0 76.1 74.0 92.0 83.0 30.0 76.0 53.0 67.0 42.6 54.8 13.0 35.0 24.0 43.3 44.7 44.0 51.3 58.9 55.1
 Table 3: Detailed performance breakdown (S/U/M) across all seven datasets.
@@
 Checkpoint Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R
-Deepseek-V3.2-R 87.8 0.00 87.8 88.8 8.30 97.1 90.0 6.00 96.0 78.0 9.00 87.0 99.5 0.00 99.5 98.0 0.00 98.0 74.4 0.80 75.2 88.1 3.44 91.5
-GPT-5.1-Low 32.0 1.00 33.0 45.1 2.08 47.2 68.0 8.00 76.0 9.0 1.00 10.0 97.5 0.00 97.5 90.0 0.00 90.0 55.3 7.90 63.2 56.7 2.85 59.6
-Gemini-3.0-Pro 94.0 1.00 95.0 99.0 0.00 99.0 95.0 0.00 95.0 99.0 0.00 99.0 96.8 0.00 96.8 93.5 0.00 93.5 64.5 0.00 64.5 91.7 0.14 91.8
-Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 38.9 0.25 39.2 46.6 0.00 46.6 41.9 5.86 47.7
-+ TruthRL 91.0 0.50 91.5 59.2 34.15 93.4 69.5 16.00 85.5 26.5 18.50 45.0 83.6 2.50 86.1 43.8 11.00 54.8 46.5 0.00 46.5 60.0 11.81 71.8
-+ w/o UnsData w/o P 87.5 0.00 87.5 58.8 28.85 87.7 65.5 15.50 81.0 33.0 11.00 44.0 82.6 0.00 82.6 47.5 0.00 47.5 51.5 0.00 51.5 60.9 7.91 68.8
-+ w/o UnsData w/ P 51.5 0.00 51.5 53.5 21.09 74.6 57.0 7.00 64.0 30.3 2.25 32.6 89.0 0.00 89.0 64.1 0.13 64.2 43.7 0.00 43.7 55.6 4.35 60.0
-+ w/ UnsData w/o P 89.5 1.50 91.0 67.8 19.00 86.8 76.5 10.00 86.5 66.0 13.00 79.0 95.4 1.00 96.4 64.8 4.00 68.8 64.7 0.42 65.1 75.0 7.00 82.0
-+ w/ UnsData w/ P 92.5 1.00 93.5 65.7 19.75 85.5 75.5 15.50 91.0 59.8 7.50 67.3 94.8 0.00 94.8 66.6 1.00 67.6 61.1 0.00 61.1 73.7 6.39 80.1
-+ Fixed-$\tau$ Final 93.3 0.00 93.3 69.6 18.49 88.1 83.3 12.00 95.3 74.5 0.00 74.5 98.2 0.00 98.2 74.2 0.00 74.2 66.3 0.00 66.3 79.9 4.36 84.3
-+ w/ UnsData w/ P Final 96.8 0.00 96.8 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.7 0.00 97.7 77.1 0.00 77.1 65.9 0.00 65.9 80.1 6.87 87.0
-Qwen3-1.7B-Instruct 50.5 0.00 50.5 25.5 8.85 34.3 23.5 12.00 35.5 14.5 0.50 15.0 42.8 0.00 42.8 40.0 0.00 40.0 35.9 0.00 35.9 33.2 3.05 36.3
-+ TruthRL 78.0 2.00 80.0 34.4 29.58 64.0 37.5 39.50 77.0 14.5 16.50 31.0 42.0 12.78 54.8 36.8 24.00 60.8 39.7 5.68 45.4 40.4 18.58 59.0
-+ w/o UnsData w/o P 82.5 0.50 83.0 42.5 29.17 71.7 49.0 30.00 79.0 29.0 6.50 35.5 47.1 0.25 47.3 39.8 3.50 43.3 35.6 4.17 39.8 46.5 10.58 57.1
-+ w/o UnsData w/ P 43.5 0.25 43.8 16.5 19.01 35.5 19.8 8.00 27.8 7.3 5.25 12.5 0.0 0.00 0.0 0.0 0.00 0.0 20.7 0.00 20.7 15.4 4.64 20.1
-+ w/ UnsData w/o P 83.5 1.00 84.5 58.5 22.88 81.4 59.5 19.50 79.0 41.5 2.50 44.0 47.9 1.50 49.4 43.0 0.75 43.8 44.0 2.52 46.5 54.0 7.24 61.2
-+ w/ UnsData w/ P 89.3 0.00 89.3 51.6 27.77 79.4 55.0 19.50 74.5 35.0 2.50 37.5 45.4 0.00 45.4 37.2 4.50 41.7 38.6 1.46 40.1 50.3 7.96 58.3
-+ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 44.1 3.17 47.3 52.7 9.52 62.2
+Deepseek-V3.2-R 87.8 0.00 87.8 88.8 8.30 97.1 90.0 6.00 96.0 78.0 9.00 87.0 99.5 0.00 99.5 98.0 0.00 98.0 62.7 0.80 63.5 86.5 3.44 90.0
+GPT-5.1-Low 32.0 1.00 33.0 45.1 2.08 47.2 68.0 8.00 76.0 9.0 1.00 10.0 97.5 0.00 97.5 90.0 0.00 90.0 52.1 7.90 60.0 56.2 2.85 59.1
+Gemini-3.0-Pro 94.0 1.00 95.0 99.0 0.00 99.0 95.0 0.00 95.0 99.0 0.00 99.0 96.8 0.00 96.8 93.5 0.00 93.5 58.1 0.00 58.1 90.8 0.14 90.9
+Qwen3-4B-Instruct 49.0 0.00 49.0 47.3 24.74 72.0 46.8 12.00 58.8 20.5 4.00 24.5 44.1 0.00 44.1 38.9 0.25 39.2 41.4 0.00 41.4 41.1 5.86 47.0
++ TruthRL 91.0 0.50 91.5 59.2 34.15 93.4 69.5 16.00 85.5 26.5 18.50 45.0 83.6 2.50 86.1 43.8 11.00 54.8 43.3 0.00 43.3 59.6 11.81 71.4
++ w/o UnsData w/o P 87.5 0.00 87.5 58.8 28.85 87.7 65.5 15.50 81.0 33.0 11.00 44.0 82.6 0.00 82.6 47.5 0.00 47.5 46.2 0.00 46.2 60.2 7.91 68.1
++ w/o UnsData w/ P 51.5 0.00 51.5 53.5 21.09 74.6 57.0 7.00 64.0 30.3 2.25 32.6 89.0 0.00 89.0 64.1 0.13 64.2 38.2 0.00 38.2 54.8 4.35 59.2
++ w/ UnsData w/o P 89.5 1.50 91.0 67.8 19.00 86.8 76.5 10.00 86.5 66.0 13.00 79.0 95.4 1.00 96.4 64.8 4.00 68.8 49.0 0.42 49.4 72.7 7.00 79.7
++ w/ UnsData w/ P 92.5 1.00 93.5 65.7 19.75 85.5 75.5 15.50 91.0 59.8 7.50 67.3 94.8 0.00 94.8 66.6 1.00 67.6 49.4 0.00 49.4 72.0 6.39 78.4
++ Fixed-$\tau$ Final 93.3 0.00 93.3 69.6 18.49 88.1 83.3 12.00 95.3 74.5 0.00 74.5 98.2 0.00 98.2 74.2 0.00 74.2 55.5 0.00 55.5 78.4 4.36 82.8
++ w/ UnsData w/ P Final 96.8 0.00 96.8 67.8 26.82 94.6 75.8 19.00 94.8 79.0 2.25 81.3 97.7 0.00 97.7 77.1 0.00 77.1 55.0 0.00 55.0 78.6 6.87 85.5
+Qwen3-1.7B-Instruct 50.5 0.00 50.5 25.5 8.85 34.3 23.5 12.00 35.5 14.5 0.50 15.0 42.8 0.00 42.8 40.0 0.00 40.0 29.8 0.00 29.8 32.4 3.05 35.5
++ TruthRL 78.0 2.00 80.0 34.4 29.58 64.0 37.5 39.50 77.0 14.5 16.50 31.0 42.0 12.78 54.8 36.8 24.00 60.8 29.6 5.68 35.3 39.0 18.58 57.6
++ w/o UnsData w/o P 82.5 0.50 83.0 42.5 29.17 71.7 49.0 30.00 79.0 29.0 6.50 35.5 47.1 0.25 47.3 39.8 3.50 43.3 28.7 4.17 32.9 45.5 10.58 56.1
++ w/o UnsData w/ P 43.5 0.25 43.8 16.5 19.01 35.5 19.8 8.00 27.8 7.3 5.25 12.5 0.0 0.00 0.0 0.0 0.00 0.0 20.7 0.00 20.7 15.4 4.64 20.0
++ w/ UnsData w/o P 83.5 1.00 84.5 58.5 22.88 81.4 59.5 19.50 79.0 41.5 2.50 44.0 47.9 1.50 49.4 43.0 0.75 43.8 39.3 2.52 41.8 53.3 7.24 60.5
++ w/ UnsData w/ P 89.3 0.00 89.3 51.6 27.77 79.4 55.0 19.50 74.5 35.0 2.50 37.5 45.4 0.00 45.4 37.2 4.50 41.7 31.0 1.46 32.5 49.2 7.96 57.2
++ w/ UnsData w/ P Final 92.0 0.50 92.5 54.4 32.19 86.6 56.0 21.50 77.5 35.0 2.00 37.0 47.6 0.55 48.2 39.5 6.75 46.3 32.1 3.17 35.3 50.9 9.52 60.4
+Llama3-8B-Distill 38.0 22.00 60.0 14.6 43.75 58.4 7.0 34.00 41.0 31.0 56.00 87.0 22.9 34.19 57.1 26.0 44.50 70.5 37.2 6.78 44.0 25.2 34.46 59.7
++ w/o UnsData w/o P 30.0 0.00 30.0 20.8 5.21 26.0 25.0 4.00 29.0 4.0 6.00 10.0 37.2 1.00 38.2 8.0 4.00 12.0 28.1 0.83 28.9 21.9 3.01 24.9
++ w/ UnsData w/ P 35.0 0.00 35.0 68.8 16.67 85.5 69.0 3.00 72.0 44.0 19.00 63.0 31.6 3.00 34.6 13.5 8.50 22.0 40.5 0.83 41.3 43.2 7.29 50.5
++ w/ UnsData w/ P Final 51.0 2.00 53.0 76.1 10.42 86.5 83.0 4.00 87.0 53.0 19.00 72.0 54.8 4.05 58.9 24.0 17.00 41.0 44.0 3.34 47.3 55.1 8.54 63.7
 Table 4: Alternative metric analysis showing Correctness (Corr), Rejection Rate (Rej), and their sum (C+R).
@@
 Finally, we present results for the fully converged models trained for 320 steps, including a variant using a static confidence threshold (Fixed-tau Final) and our fully converged model trained with the complete UnsolvableRL method (w/ UnsData w/ P Final).
+We also extended our evaluation to the Llama3-8B architecture to verify the generalizability of our method.
+As shown in the tables, training exclusively on solvable data (w/o UnsData w/o P) yields a dramatic improvement in solvable accuracy compared to the baseline ($S:15.9\%\to 39.6\%$).
+However, this comes at a catastrophic cost: the model’s ability to detect unsolvable problems collapses ($U:34.6\%\to 4.1\%$), leading to a net decline in the overall Mean score ($M:25.2\%\to 21.9\%$).
+In contrast, our proposed method (w/ UnsData w/ P) successfully resolves this conflict.
+At step 120, it maintains superior solvable accuracy ($S=40.1\%$) while simultaneously robustifying unsolvable detection ($U=46.3\%$), nearly doubling the overall Mean score to 43.2%.
+Furthermore, extending the training to approximately 400 steps (Final) demonstrates robust scalability: the model achieves further gains with Solvable Accuracy reaching 51.3% and Unsolvable Detection rising to 58.9%, resulting in a comprehensive Mean score of 55.1%.
+This confirms that UnsolvableRL effectively prevents capability collapse and enhances reasoning rigor across different model families.
 A.2 Out-of-Distribution (OOD) Generalization
@@
 Table 5: OOD performance. S: Solvable, U: Unsolvable, M: Mean accuracy (%).
-A.3 Theoretical Justification for False Unsolvability Penalty
+A.3 Case Study
+To qualitatively demonstrate the effectiveness of our alignment strategy, we present three representative cases generated by DeepSeek-V3.2-R.
+These examples illustrate the model’s ability to navigate complex reasoning paths: detecting latent contradictions in mathematics (Case 1), identifying topological impossibility in logic puzzles (Case 2), and maintaining flexible reasoning for solvable problems (Case 3).
+Case 1: Latent Unsolvability (Number Theory)
+Problem Input.
+Let $N$ be the greatest four-digit positive integer with the property that whenever one of its digits is changed to $1$, the resulting number is divisible by $7$.
+Suppose that $N\equiv Q+R\pmod{7}$, where $Q$ and $R$ are the quotient and remainder when $N$ is divided by $1000$.
+Find $Q+R$.
+Original Model Response.
+“The digit-change property requires that for any digit of $N$, replacing that digit with 1 results in a number divisible by 7.
+Under the strict interpretation… the only solutions are $N=5624$ and $N=5694$, with the latter being the greatest.
+However, for $N=5694$, computing $Q=5$ and $R=694$ gives $Q+R=699$.
+Checking the congruence $N\equiv Q+R\pmod{7}$ yields $5694\equiv 3\pmod{7}$ and $699\equiv 6\pmod{7}$, which are not equal.
+… Thus, the greatest $N$ satisfying the digit-change property does not satisfy the additional condition.
+Therefore, no four-digit $N$ meets all the requirements.”
+$\boxed{\text{unsolvable}}$
+Reasoning Analysis.
+This case exemplifies "Verify-then-Conclude". The contradiction is not explicit in the problem statement but lies deep within the solution space.
+  • Deep Execution: The model does not reject the problem early. It successfully solves the core number theory challenge, identifying the unique candidate $N=5694$.
+  • Constraint Verification: Instead of hallucinating a match, it rigorously computes the modular arithmetic ($3\neq 6$) and identifies the conflict between the derived candidate and the final constraint.
+Case 2: Topological Impossibility (Hitori)
+Problem Input.
+Solve the following $4\times 4$ Hitori puzzle. Rules: No duplicates in rows/cols; Black cells cannot touch; White cells must be connected.
+4 2 3 3
+2 1 4 2
+4 3 3 2
+4 1 2 2
+Original Model Response.
+“The puzzle is unsolvable due to the isolation of cell (0,3) in all valid configurations… Specifically, the triplet of 2s in column 3 forces the outer cells (1,3) and (3,3) to be black to avoid adjacent shading.
+This forces (0,3) to be white.
+However, row 0 has duplicate 3s at (0,2) and (0,3), forcing (0,2) to be black.
+This leaves the white cell (0,3) surrounded by black cells (0,2) and (1,3), disconnecting it from the rest of the board.”
+$\boxed{\text{unsolvable}}$
+Reasoning Analysis.
+This demonstrates the model’s grasp of Global Constraints.
+  • Logical Deduction: The model correctly deduces that to resolve duplicates in Col 3 (values: 3,2,2,2) and Row 0 (values: 4,2,3,3), specific neighbors of the corner cell (0,3) must be shaded.
+  • Connectivity Check: It recognizes that satisfying local uniqueness constraints inevitably violates the global connectivity rule (isolating the white cell at 0,3).
+This shows the model isn’t just counting numbers but understanding the topological structure of the puzzle.
+Case 3: Flexible Reasoning (Game24)
+Problem Input.
+Use the numbers $\{4,4,6,12\}$ to compute 24.
+Original Model Response.
+“To find an arithmetic expression that equals 24… One straightforward approach is to obtain a 4 from the numbers 12, 4, and 4, and then multiply the result by 6.
+1.
+Start with 12.
+Subtract 4: 124=812-4=8.
+2.
+Subtract the second 4: 84=48-4=4.
+3.
+Multiply the result by 6: $4\times 6=24$.
+This gives the expression: $(12-4-4)\times 6$.”
+(1244)6(12-4-4)*6
+Reasoning Analysis.
+  • No False Refusal: Despite the rigorous training on unsolvable data, the model correctly identifies a valid path for a solvable query.
+  • Flexibility: The model finds a subtraction-based solution different from the standard division-based answer ($4\times(12/(6-4))$), showing that the alignment process has not constrained its exploration capability.
+Summary.
+Together, these cases validate that DeepSeek-R aligned with UnsolvableRL does not rely on simple heuristics to reject questions.
+It performs comprehensive reasoning—whether calculating modular congruences, checking graph connectivity, or exploring arithmetic combinations—before making a solvability determination.
+Game24 HamCycle HamPath Hitori Maze(E) Maze(H) AIME24-25 Overall
+Checkpoint S U M S U M S U M S U M S U M S U M S U M $\mathrm{S}$ $\mathrm{U}$ $\mathrm{M}$
+4B UnsolvableRL (P=-0.5, S40) 79.0 93.0 86.0 29.2 98.0 63.6 48.0 98.0 73.0 37.0 41.0 39.0 53.0 91.0 72.0 13.0 78.0 45.5 70.0 31.9 51.0 47.0 75.8 61.4
+4B UnsolvableRL (P=-1.0, S40) 74.0 90.0 82.0 35.4 98.0 66.7 34.0 98.0 66.0 50.0 36.0 43.0 48.0 74.5 61.2 22.0 78.0 50.0 68.3 29.7 49.0 47.4 72.0 59.7
+4B UnsolvableRL (P=-2.0, S40) 80.0 90.0 85.0 29.2 98.0 63.6 40.0 98.0 69.0 42.0 22.0 32.0 61.0 70.7 65.8 15.5 76.5 46.0 67.5 29.7 48.6 47.9 69.4 58.6
+Table 6: Detailed performance breakdown (S/U/M) across all seven datasets for different Penalty values at 40 steps.
+A.4 Theoretical Justification and Empirical Ablation for False Unsolvability Penalty
+A.5 Theoretical Justification and Empirical Ablation for False Unsolvability Penalty
 We explicitly motivate the penalty value $\rho=-0.5$ via a decision-theoretic analysis.
@@
 This “Threshold Expansion” effectively encourages the model to attempt reasoning even under moderate uncertainty, preventing the collapse of reasoning capabilities.
-A.4 Theoretical Analysis of Capability-Refusal Dynamics
-In this section, we derive the mathematical relationship between the model’s reasoning capability and its tendency to refuse, explaining why a dynamic threshold is essential.
-The model should choose to refuse only if the reward for refusal exceeds the expected reward of answering:
-$R_{\text{cal}}>\mathbb{E}[R_{\text{ans}}].$ (6)
-Note that the expected reward for answering is always non-negative ($\mathbb{E}[R_{\text{ans}}]\geq 0$).
-Failure of Fixed Threshold.
-If we were to use a fixed threshold (e.g., $\tau=0.5$), as the model’s reasoning capability improves during training, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$).
-Once this occurs, the refusal reward becomes negative ($R_{\text{cal}}=\lambda(\tau-\beta)<0$).
-Substituting this into Inequality 6, the condition becomes $\text{Negative Value}>\mathbb{E}[R_{\text{ans}}]$.
-Since $\mathbb{E}[R_{\text{ans}}]\geq 0$, this inequality is impossible to satisfy.
-Consequently, for well-performed batches, the model is incentivized to suppress refusal entirely to avoid penalties, effectively forcing the model to hallucinate answers on difficult problems rather than refusing.
-Our progressive schedule prevents this collapse by ensuring $\tau$ grows dynamically to maintain $\tau\gtrsim\beta$, keeping the refusal incentive positive for uncertain queries.
-A.5 Logic Puzzle Generation Algorithms
-A.5.1 Game24 Generation
-The Game24 task involves combining a set of integers using arithmetic operations (+,,×,÷)(+,-$\times,\div$) to equal 24.
-To ensure absolute correctness, particularly for unsolvable instances, we employ an exhaustive search over all possible binary expression trees.
-Crucially, we utilize exact rational arithmetic (e.g., Python’s Fraction) rather than floating-point calculations to prevent precision errors that could misclassify instances (e.g., $23.99999\neq 24$).
-For unsolvable instances, we use a rejection sampling strategy: we randomly sample integer sets and retain only those for which the exhaustive solver proves the solution set is empty.
+To empirically validate this theoretical derivation, we conducted an ablation study on the penalty magnitude using the Qwen3-4B model at 40 training steps, as detailed in Table 6.
+We compared the default $\rho=-0.5$ against stricter penalties ($\rho=-1.0$ and $\rho=-2.0$).
+The results indicate that while the method is generally robust, continuing to decrease $\rho$ (making the penalty more severe) is counterproductive.
+Specifically, decreasing $\rho$ to 2.0-2.0 causes the Unsolvable Detection Rate (Overall U) to drop significantly from 75.8% to 69.4%, as the model becomes overly hesitant to declare unsolvability.
+Crucially, this sacrifice in detection yields only a marginal gain in Solvable Accuracy (Overall S increases from 47.0% to 47.9%), resulting in a decline in the global Mean score from 61.4% to 58.6%.
+Consequently, $\rho=-0.5$ proves to be the optimal value, striking the best balance between maintaining reasoning rigor and preserving the courage to reject unsolvable queries.
+To ensure reliability, we report the average results across two independent training runs, evaluated using mean@2.
 Algorithm 1 Game24 Generation and Verification
@@
 16:until Valid instance found
-A.5.2 Hamiltonian Cycle and Path Generation
+Game24 HamCycle HamPath Hitori Maze(E) Maze(H) AIME24-25 Overall
+Checkpoint Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R Corr Rej C+R
+4B UnsolvableRL ($\tau$=0.2, S40) 90.5 3.00 93.5 63.7 23.21 86.9 72.5 10.50 83.0 32.5 6.50 39.0 69.0 8.75 77.7 46.5 21.00 67.5 49.5 0.00 49.5 60.6 10.42 71.0
+4B UnsolvableRL ($\tau$=0.4, S40) 86.0 1.50 87.5 63.6 24.48 88.1 73.0 5.50 78.5 39.0 10.00 49.0 72.0 4.50 76.5 45.5 18.50 64.0 51.0 0.21 51.2 61.4 9.17 70.6
+4B UnsolvableRL ($\tau$=0.6, S40) 86.0 3.00 89.0 58.2 31.19 89.4 66.0 13.50 79.5 39.5 8.00 47.5 71.3 6.75 78.1 47.8 26.50 74.3 48.0 0.83 48.8 59.5 12.82 72.3
+4B UnsolvableRL ($\tau$=0.8, S40) 78.0 3.00 81.0 67.3 30.69 98.0 70.5 13.50 84.0 36.0 8.00 44.0 70.7 6.75 77.5 49.3 26.25 75.6 46.4 0.84 47.2 59.7 12.72 72.4
+4B UnsolvableRL ($\tau$=1.0, S40) 85.5 3.50 89.0 59.7 35.19 94.9 65.5 21.50 87.0 26.5 24.50 51.0 78.8 2.75 81.6 47.0 26.25 73.3 49.3 0.42 49.7 58.9 16.30 75.2
+Table 7: Detailed performance breakdown (Correctness/Rejection Rate/C+R) across all seven datasets for different $\tau$ values at 40 steps.
+A.6 Theoretical Justification and Empirical Ablation for Capability-Refusal Dynamics
+In this section, we derive the mathematical relationship between the model’s reasoning capability and its tendency to refuse, explaining why a dynamic threshold is essential.
+The model should choose to refuse only if the reward for refusal exceeds the expected reward of answering:
+$R_{\text{cal}}>\mathbb{E}[R_{\text{ans}}].$ (6)
+Note that the expected reward for answering is always non-negative ($\mathbb{E}[R_{\text{ans}}]\geq 0$).
+Failure of Fixed Threshold.
+If we were to use a fixed threshold (e.g., $\tau=0.5$), as the model’s reasoning capability improves during training, the batch accuracy $\beta$ will likely surpass $\tau$ (i.e., $\beta>\tau$).
+Once this occurs, the refusal reward becomes negative ($R_{\text{cal}}=\lambda(\tau-\beta)<0$).
+Substituting this into Inequality 6, the condition becomes $\text{Negative Value}>\mathbb{E}[R_{\text{ans}}]$.
+Since $\mathbb{E}[R_{\text{ans}}]\geq 0$, this inequality is impossible to satisfy.
+Consequently, for well-performed batches, the model is incentivized to suppress refusal entirely to avoid penalties, effectively forcing the model to hallucinate answers on difficult problems rather than refusing.
+Our progressive schedule prevents this collapse by ensuring $\tau$ grows dynamically to maintain $\tau\gtrsim\beta$, keeping the refusal incentive positive for uncertain queries.
+Specifically, in our experiments, the Fixed-$\tau$ baseline uses a constant $\tau=0.4$, while our dynamic schedule linearly increases $\tau$ from $0.4$ to $1.0$ over 400 training steps.
+To further validate our threshold selection, we conducted a simple ablation study on $\tau$.
+As shown in Table 7, we found that when training for only 40 steps, setting $\tau=0.4$ generally yields strong Accuracy, while further decreasing $\tau$ has minimal impact on performance.
+Based on these observations, we adopted $\tau=0.4$ as the initial value for our main experiments.
+A.7 Logic Puzzle Generation Algorithms
+A.7.1 Game24 Generation
+The Game24 task involves combining a set of integers using arithmetic operations (+,,×,÷)(+,-$\times,\div$) to equal 24.
+To ensure absolute correctness, particularly for unsolvable instances, we employ an exhaustive search over all possible binary expression trees.
+Crucially, we utilize exact rational arithmetic (e.g., Python’s Fraction) rather than floating-point calculations to prevent precision errors that could misclassify instances (e.g., $23.99999\neq 24$).
+For unsolvable instances, we use a rejection sampling strategy: we randomly sample integer sets and retain only those for which the exhaustive solver proves the solution set is empty.
+A.7.2 Hamiltonian Cycle and Path Generation
 The Hamiltonian Cycle problem asks whether a graph contains a closed loop visiting every vertex exactly once, while the Hamiltonian Path problem asks for a linear path visiting every vertex.
@@
 19:return $G$
-A.5.3 Hitori Generation
+A.7.3 Hitori Generation
 Hitori is a logic puzzle played on an $N\times N$ grid.
@@
 19:until Valid instance found
-A.5.4 Maze Navigation Generation
-The Maze Navigation task requires finding a valid path from a start position (S) to an end position (E) in a grid-based maze, where movement is restricted by walls.
-Verification Logic.
-We employ Breadth-First Search (BFS) to determine solvability.
-BFS guarantees finding the shortest path if one exists.
-A maze is Solvable if a path from S to E exists, and Unsolvable if the set of reachable cells from S does not contain E.
-Construction Pipeline.
-We first generate a complex solvable maze using a randomized Depth-First Search (DFS) algorithm, which ensures a spanning tree structure.
-To increase complexity, we selectively remove walls to introduce cycles and multiple potential paths, while validating that the solution remains unique or manageable using BFS.
-To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement.
-We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction, a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths).
-We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS.
-If the maze remains solvable (e.g., via a detour), we discard the instance or iteratively block remaining paths until connectivity is severed.
-This approach ensures that unsolvable mazes are structurally similar to solvable ones, often requiring the model to explore deep into the maze before realizing the path is blocked, thereby penalizing premature "give up" or hallucinated paths.
 Algorithm 4 Maze Generation with Strategic Blockage
@@
 15:until Valid instance found
-A.6 Hyperparameters
-We implement our training pipeline using the verl (HybridFlow) framework and vllm engine (sheng2025hybridflow; kwon2023efficient), which supports efficient reinforcement learning.
-The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 6.
-We set the group size $G$ (denoted as rollout.n in verl) to 12.
-Notably, we explicitly set use_kl_loss=False, meaning the KL divergence penalty term was disabled during optimization, relying instead on the probability ratio clipping to maintain policy stability.
-We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints.
-Specifically, the 1.7B model utilizes a larger batch size ($32\times 12$) compared to the 4B model ($14\times 12$).
-This adjustment balances the computational overhead, resulting in comparable training speeds for both architectures; typically, completing 120 training steps requires approximately one day on our 8 $\times$ NVIDIA A100 GPU cluster.
 Table 8: Hyperparameter settings used in our experiments.
@@
    Max Response Length 32,000
+A.7.4 Maze Navigation Generation
+The Maze Navigation task requires finding a valid path from a start position (S) to an end position (E) in a grid-based maze, where movement is restricted by walls.
+Verification Logic.
+We employ Breadth-First Search (BFS) to determine solvability.
+BFS guarantees finding the shortest path if one exists.
+A maze is Solvable if a path from S to E exists, and Unsolvable if the set of reachable cells from S does not contain E.
+Construction Pipeline.
+We first generate a complex solvable maze using a randomized Depth-First Search (DFS) algorithm, which ensures a spanning tree structure.
+To increase complexity, we selectively remove walls to introduce cycles and multiple potential paths, while validating that the solution remains unique or manageable using BFS.
+To construct Unsolvable Instances, we employ a Strategic Blockage method rather than random wall placement.
+We analyze the solvable maze to identify all valid paths from S to E, and then select a Critical Junction, a coordinate $(x,y)$ that lies on a valid path (or is shared by multiple paths).
+We place an obstacle (wall) at $(x,y)$ and rigorously verify unsolvability using BFS.
+If the maze remains solvable (e.g., via a detour), we discard the instance or iteratively block remaining paths until connectivity is severed.
+This approach ensures that unsolvable mazes are structurally similar to solvable ones, often requiring the model to explore deep into the maze before realizing the path is blocked, thereby penalizing premature "give up" or hallucinated paths.
+A.8 Hyperparameters and Training Details
+Figure 7: Average response token length trend during training.
+The solid lines represent the moving average, and the shaded regions indicate the volatility ($\pm 1\sigma$).
+Both models show a tendency to increase reasoning length as training progresses.
+We implement our training pipeline using the verl (HybridFlow) framework and vllm engine (Sheng et al., 2025; Kwon et al., 2023), which supports efficient reinforcement learning.
+The specific hyperparameters used for Group Relative Policy Optimization (GRPO) are detailed in Table 8.
+We set the group size $G$ (denoted as rollout.n in verl) to 12.
+Notably, we explicitly set use_kl_loss=False, meaning the KL divergence penalty term was disabled during optimization, relying instead on the probability ratio clipping to maintain policy stability.
+We differentiate the batch size for Qwen3-1.7B and Qwen3-4B to accommodate their respective memory footprints.
+Specifically, the 1.7B model utilizes a larger batch size ($32\times 12$) compared to the 4B model ($14\times 12$).
+This adjustment balances the computational overhead, resulting in comparable training speeds for both architectures; typically, completing 120 training steps requires approximately one day on our 8 $\times$ NVIDIA A100 GPU cluster.
+Constrained by GPU resources, we conducted a single training run for the main experiments and report results using mean@8 for AIME and mean@4 for the remaining benchmarks.
+Throughout the training process, we observed a stable increase in both Solvable Accuracy and Unsolvable Detection Rate.
+To further analyze the training dynamics, Figure 7 illustrates the evolution of the average response token length throughout the training process.
+We observe a general upward trend for both models, indicating that the RL alignment encourages the generation of more comprehensive reasoning chains (longer CoT) to distinguish solvability.
+The Qwen3-4B model exhibits higher volatility (indicated by the wider shaded area) compared to the 1.7B model, suggesting more aggressive exploration within the solution space.
+[…]
+We will release our dataset, data construction pipeline, and training code upon acceptance.